### v2

# Contextualized Messages Boost Graph Representations

Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This success has prompted several studies to explore the representational capability of GNNs based on the graph isomorphism task. These works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few theoretical works study GNNs with uncountable node feature representation. This paper presents a novel perspective on the representational capability of GNNs across all levels – node-level, neighborhood-level, and graph-level – when the space of node feature representation is uncountable. Specifically, it relaxes the injective requirement in previous works by employing an implicit pseudometric distance on the space of input to create a soft-injective function. This allows distinct inputs to produce similar outputs only if the pseudometric deems the inputs to be sufficiently similar on some representation, which is often useful in practice. As a consequence, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes non-linear and contextualized transformation of neighborhood feature representations is proposed. A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution in context, establishing SIR-GCN as a generalization of classical GNN methodologies. Experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others [12, 19]. GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation [14, 47, 48]. This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph [27, 32, 47, 50].
Among the most widely used GNNs in literature include the graph convolution network (GCN) [25], graph sample and aggregate (GraphSAGE) [15], graph attention network (GAT) [7, 40], and graph isomorphism network (GIN) [47] which largely fall under the message-passing neural networks (MPNNs) [14] framework. These models have gained popularity due to their simplicity and remarkable performance across various applications [12, 17, 19, 22, 24, 28]. Improvements are also constantly being proposed to achieve state-of-the-art performance [4, 6, 21, 31, 38, 44, 49].
Notably, these advances are mainly driven by heuristics and empirical results. Nonetheless, several studies have also begun exploring the representational capability of GNNs [2, 4, 5, 11, 13, 34]. Most of these works analyzed GNNs in relation to the graph isomorphism task. Xu et al. [47] was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test [46]. Subsequent works build upon their results by considering extensions to the original 1-WL test. However, these results only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. [11] proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
This paper presents a simple yet novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable. The key idea is to define an implicit pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This idea is explored across all levels – node-level, neighborhood-level, and graph-level. Based on the results, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the non-linear and contextualized transformation of neighborhood feature representations is proposed. The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the advantages of the proposed model. Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN in closing.
2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{N}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{N}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
$\bm{H_{u}}$ =$=\left\{\!\!\left\{\bm{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}$
$\bm{a_{u}}$ $=\textsc{AGG}\left(\bm{H_{u}}\right)$ (1)
$\bm{h^{*}_{u}}$ $=\textsc{COMB}\left(\bm{h_{u}},\bm{a_{u}}\right),$
where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset [47] of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
Related works
When $\mathcal{H}$ is countable, Xu et al. [47] showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
$F\left(\bm{H}\right)=\sum_{\bm{h}\in\bm{H}}f\left(\bm{h}\right)$ (2)
is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size. This result forms the theoretical basis of GIN.
Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable. In this setting, Corso et al. [11] proved that if $\bigoplus$ comprises multiple aggregators (e.g. mean, standard deviation, max, and min), the hash function
$M\left(\bm{H}\right)=\bigoplus_{\bm{h}\in\bm{H}}m\left(\bm{h}\right)$ (3)
produces a unique output for every $\bm{H}$ of bounded size. This finding laid the foundation for the principal neighborhood aggregation (PNA) [11]. Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
3 Soft-injective functions
Theorem 1 presents an alternative to injective functions when the space of node features $\mathcal{H}$ is uncountable. It considers a soft relaxation to injectivity, henceforth soft-injectivity, which is often useful in practice, especially in classification tasks.
Theorem 1.
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
Theorem 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function operating on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bounds $\delta_{2}$ and $\varepsilon_{2}$ assert the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bounds $\delta_{1}$ and $\varepsilon_{1}$ ensure $g_{u}$ maintains the relationship of elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ is then said to be soft-injective. An illustration is provided in Fig. 1.
$\delta$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(a) Embedded feature space $\mathcal{S}$.
$d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)<\varepsilon<d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(b) Input feature space $\mathcal{H}$.
Figure 1: Soft-injective feature map $g_{u}$ with pseudometric $d_{u}$.
It is worth noting that if $d_{u}$ is squared Euclidean, then Theorem 1 becomes trivial. Nevertheless, the result becomes non-trivial for other choices of $d_{u}$. Corollary 1 extends this result for multisets.
Corollary 1.
Let $\mathcal{H}$ be a non-empty set with a pseudometric $D$ on bounded, equinumerous multisets of $\mathcal{H}$ defined as
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (5)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
where
$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h}).$ (7)
3.1 Soft-isomorphic relational graph convolution network
Corollary 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn. 7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Similarly, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the message function [14] of the aggregation strategy that transforms the individual neighborhood features. Meanwhile, the distance function $D_{u}$ may be interpreted as a kernel distance [23].
Definition 1 (Collision).
Let $G$ be a function. If $G\left(\bm{H}^{(1)}\right)=G\left(\bm{H}^{(2)}\right)$ and $\bm{H}^{(1)}\neq\bm{H}^{(2)}$, a collision is said to have occurred.
To illustrate the utility of $D_{u}$, suppose node $u$ has two neighbors $v_{1}$, $v_{2}$. If $d_{u}$ is the squared Euclidean distance, then a corresponding message function $g_{u}$ is linear. Fig. 2(a) presents the contour plot of the corresponding hash function $G_{u}$, highlighting potential issues arising from hash collisions. Specifically, consider two multisets $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$. If $\bm{h_{v_{1}}}^{(1)}+\bm{h_{v_{2}}}^{(1)}=\bm{h_{v_{1}}}^{(2)}+\bm{h_{v_{2}}}^{(2)}$, then a hash collision occurs and $G_{u}$ produces identical aggregated neighborhood features even if $\bm{H_{u}}^{(1)}$ and $\bm{H_{u}}^{(2)}$ are fundamentally dissimilar on some representation for a given task.
[…]
(a) $g_{u}(\bm{h})=\bm{h}$.
[…]
(b) $g_{u}(\bm{h})=-\bm{h}^{2}$.
[…]
(c) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h})$.
[…]
(d) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h}+\bm{1})$.
Figure 2: Hash functions $G_{u}$ under different message functions $g_{u}$.
In general, hash collisions may occur when aggregating neighborhood features due to $\mathcal{H}$ being uncountable. Within the current framework, one may simply encode knowledge about node features into the pseudometric $D_{u}$. As a result, only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled. This also corresponds to a more complex and non-linear message function $g_{u}$. To illustrate, if node features represent a zero-mean score, $d_{u}$ may be defined as the squared Euclidean distance of the squared score. A corresponding hash function $G_{u}$ in Fig. 2(b) may then be used to detect potentially anomalous neighborhoods since hash collisions are meaningful for this task.
It is worth noting, however, that Corollary 1 holds for every node $u\in\mathcal{N}$ independently. Hence, different nodes may correspond to different $D_{u}$ and $G_{u}$. For simplicity, one may consider only a single pseudometric compactly defined as
$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})$ (8)
with a corresponding soft-injective hash function
$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right)$ (9)
for every node $u\in\mathcal{N}$. This approach preserves the interpretation of $G$ as an aggregation or hash function with an underlying pseudometric distance $D$ that guides and controls hash collisions. The integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function guiding how features of the key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$ are to be embedded and transformed based on the features of the query (center) node $\bm{h_{u}}$. This provides additional context to hash collisions and makes the message function and pseudometric anisotropic [12] or adaptive with respect to the query node. Figs. 2(c) and 2(d) demonstrate this idea where the introduction of a bias term, assuming a function of $\bm{h_{u}}$, shifts the contour plot and produces different aggregated neighborhood features for identical neighborhood features. Nevertheless, one may opt to inject stochasticity into the node features to distinguish between nodes with identical features and neighborhood features and imitate having distinct hash functions $G_{u}$ for every node $u$ with high probability [34].
For a graph representation learning problem, the relational message function $g$ may be modeled as a two-layered multi-layer perceptron (MLP), with an implied pseudometric, following the universal approximation theorem [16] to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (10)
where $\sigma$ is a non-linear activation function, $\bm{W_{Q}},\bm{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\bm{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$. Leveraging linearity, the model has a computational complexity of
$\mathcal{O}\left(\left|\mathcal{N}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
which is comparable to GAT. In practice, $\sigma$ may also be replaced with a deep MLP if modeling $g$ as a shallow two-layer MLP is infeasible. Moreover, since $\bm{h_{u}}$ is already encoded in the aggregation strategy, the combination strategy may simply be an activation function.
In essence, the proposed SIR-GCN is an instance of the MPNN framework where, unlike most MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic [7] transformation of the neighborhood features to obtain contextualized messages.
3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for every graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn. 7. While this result holds for every graph $\mathcal{G}$ independently, one may simply assume a single pseudometric $D$ with a corresponding soft-injective graph readout function $R$ for a set of graphs $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ from task $\mathcal{D}$. Nevertheless, one may opt to integrate additional information about the graph context and structure into the graph readout function $R$ to imitate having distinct $R_{\mathcal{G}}$ for every graph $\mathcal{G}$ and enhance its representational capability further. The virtual super node [14] may be used in this regard.
In practice, $R$ may also be modeled with an MLP, with an implied pseudometric, to obtain the soft-isomorphic graph readout function
$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{N}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
where $\textsc{MLP}_{R}$ represents the corresponding feature map of $R$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
4 Mathematical discussion
The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the contribution. The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the proposed model.
4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation (by extension) may be used in place of the sum aggregation. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
$\bm{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\bm{W}\bm{h_{v}}$ (13)
and
$\bm{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\bm{W}\bm{h_{v}}$ (14)
which recovers GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN. [^1] GraphSAGE with LSTM aggregation is not included in this discussion. The difference lies in the isotropic [12] nature of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
4.2 GAT
Moreover, in Brody et al. [7], the attention mechanism of GATv2 is modeled as an MLP given by
$e_{u,v}=\bm{a}^{\top}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (15)
with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\bm{W}\bm{h_{v}}$. While the model is anisotropic in nature, messages are nevertheless only linearly transformed with node $u$ only determining the degree of contribution through the scalar $e_{u,v}$. Meanwhile, SIR-GCN directly works with the unnormalized attention mechanism in Eqn. 15 and allows the features of the query node to dynamically transform messages. Specifically, if $\sigma=\textsc{LeakyReLU}$ and $\bm{W_{R}}=\bm{A}$, one obtains
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{A}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right)$ (16)
which shows Eqn. 15 becoming a contextualized message in the SIR-GCN model. Nevertheless, GAT and GATv2 may be recovered, up to a normalizing constant, with an appropriate choice of $\sigma$.
4.3 GIN
Likewise, within the proposed SIR-GCN model, one may explicitly add a residual connection in the combination strategy to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (17)
where $\textsc{MLP}_{\text{Res}}$ is a learnable residual network. If $\textsc{MLP}_{\text{Res}}(\bm{h})=(1+\epsilon)\cdot\bm{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, and $\bm{W_{R}W_{K}}=\bm{I}$, then
$\bm{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\bm{h_{u}}+\sum_{v\in\mathcal{N}(u)}\bm{h_{v}}$ (18)
is equivalent to a GIN. Hence, SIR-GCN with residual connection encompasses GIN.
4.4 PNA
Furthermore, SIR-GCN and PNA approach the problem of uncountable node features $\mathcal{H}$ differently, with the former using only a single aggregator which holds theoretically for any number of neighbors. Nevertheless, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using a linear message function $m$ which is equivalent to a linear transformation of $\bm{h_{v}}$ with a different bias term for each $u$
$m\left(\bm{h_{v}},\bm{h_{u}}\right)=\bm{W_{K}}\bm{h_{v}}+\bm{W_{Q}}\bm{h_{u}}=\bm{W_{K}}\bm{h_{v}}+\bm{b_{u}}.$ (19)
When using mean, max, or min aggregators, the query node $u$ only contributes to the aggregated neighborhood features through the bias term $\bm{b_{u}}$. Meanwhile, when using normalized moment aggregators, the query node $u$ no longer contributes to the aggregated neighborhood features. This potentially limits the expressivity of PNA since, in contrast to SIR-GCN and as suggested by Brody et al. [7], messages are only linearly and statically [7] transformed by the features of the query node.
4.5 1-WL test
Additionally, in terms of graph isomorphism representational capability, SIR-GCN is comparable to a modified 1-WL test. Suppose $w^{(l)}_{u}$ is the WL node label of node $u$ at the $l$th WL-test iteration. The modified update equation is given by
$w^{(l)}_{u}\leftarrow\text{hash}\left(\left\{\!\!\!\left\{\left[w^{(l-1)}_{v},w^{(l-1)}_{u}\right]:v\in\mathcal{N}(u)\right\}\!\!\!\right\}\right),$ (20)
where the modification lies in concatenating the label of the center node with every element of the multiset before hashing. This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as noted in the previous section. Thus, SIR-GCN inherits the theoretical capabilities (and limitations) of the 1-WL test.
Overall, SIR-GCN is demonstrated to generalize four prominent GNNs in literature – GCN, GraphSAGE, GAT, and GIN – and is thus at least as expressive as they are. Notably, SIR-GCN offers flexibility in two dimensions of GNNs: aggregation strategy and message transformation. It emphasizes summed aggregation with anisotropic and dynamic (i.e. contextualized) message transformation, making it well-suited for heterophilous tasks [8], but remains adaptable to alternative configurations.
In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator, which theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
SumMeanAnisotropicIsotropicGATGINGraphSAGEGCNSIR-GCN
Figure 3: SIR-GCN expressivity.
5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN. To ensure fairness, models not employing advanced architectural design or manually crafted features using domain knowledge are used as primary comparisons.
5.1 Synthetic datasets
DictionaryLookup
DictionaryLookup [7] consists of bipartite graphs with $2n$ nodes – $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the value of query nodes by matching their attribute with the key nodes. A sample instance is provided in Fig. 4.
Table 1: Test accuracy on DictionaryLookup.
Model $n=10$ $n=20$ $n=30$ $n=40$ $n=50$
GCN 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00
GraphSAGE 0.10 $\pm$ 0.00 0.05 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
GATv2 0.99 $\pm$ 0.03 0.88 $\pm$ 0.18 0.74 $\pm$ 0.28 0.56 $\pm$ 0.37 0.60 $\pm$ 0.40
GIN 0.78 $\pm$ 0.07 0.29 $\pm$ 0.03 0.12 $\pm$ 0.03 0.03 $\pm$ 0.00 0.02 $\pm$ 0.01
SIR-GCN 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
Figure 4: DictionaryLookup.
$A$$B$$A$$B$
Figure 5: GraphHeterophily.
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $n$. SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic nature. However, it is observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic nature. The results underscore the utility of an attentional/relational mechanism in capturing the relationship between the query and key nodes.
GraphHeterophily
GraphHeterophily is an original synthetic dataset. It consists of random directed graphs with each node assigned one of $c$ classes. The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph. A sample instance is provided in Fig. 5.
Table 2: Test mean squared error on GraphHeterophily.
Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
GCN 22749 $\pm$ 1242 50807 $\pm$ 2828 62633 $\pm$ 3491 68965 $\pm$ 3784 72986 $\pm$ 4025
GraphSAGE 22962 $\pm$ 1215 36854 $\pm$ 2330 30552 $\pm$ 1574 21886 $\pm$ 1896 16529 $\pm$ 1589
GATv2 22329 $\pm$ 1307 44972 $\pm$ 2834 49940 $\pm$ 2942 50063 $\pm$ 3407 49661 $\pm$ 3488
GIN 39.620 $\pm$ 2.060 37.193 $\pm$ 1.382 34.649 $\pm$ 1.502 32.424 $\pm$ 1.841 30.091 $\pm$ 1.429
SIR-GCN 0.001 $\pm$ 0.000 0.004 $\pm$ 0.005 1.495 $\pm$ 4.428 0.038 $\pm$ 0.068 0.089 $\pm$ 0.134
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $c$. SIR-GCN achieves an MSE loss of nearly 0 attributed to its anisotropic nature and sum aggregation. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. [47]. Meanwhile, GIN successfully retains the graph structure but fails to learn the relationship between the labels of the query (center) node and key (neighboring) nodes. The results illustrate the utility of anisotropic models even in graph property prediction tasks with countable node features.
5.2 Benchmark datasets
Benchmarking GNNs
Benchmarking GNNs [12] is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. Dwivedi et al. [12] provides more information regarding the individual datasets.
Table 3: Test performance on Benchmarking GNNs.
Model WikiCS ($\uparrow$) PATTERN ($\uparrow$) CLUSTER ($\uparrow$) MNIST ($\uparrow$) CIFAR10 ($\uparrow$) ZINC ($\downarrow$)
MLP 59.45 $\pm$ 2.33 50.52 $\pm$ 0.00 20.97 $\pm$ 0.00 95.34 $\pm$ 0.14 56.34 $\pm$ 0.18 0.706 $\pm$ 0.006
GCN 77.47 $\pm$ 0.85 85.50 $\pm$ 0.05 47.83 $\pm$ 1.51 90.12 $\pm$ 0.15 54.14 $\pm$ 0.39 0.416 $\pm$ 0.006
GraphSAGE 74.77 $\pm$ 0.95 50.52 $\pm$ 0.00 50.45 $\pm$ 0.15 97.31 $\pm$ 0.10 65.77 $\pm$ 0.31 0.468 $\pm$ 0.003
GAT 76.91 $\pm$ 0.82 75.82 $\pm$ 1.82 57.73 $\pm$ 0.32 95.54 $\pm$ 0.21 64.22 $\pm$ 0.46 0.475 $\pm$ 0.007
GIN 75.86 $\pm$ 0.58 85.59 $\pm$ 0.01 58.38 $\pm$ 0.24 96.49 $\pm$ 0.25 55.26 $\pm$ 1.53 0.387 $\pm$ 0.015
GatedGCN - 84.48 $\pm$ 0.12 60.40 $\pm$ 0.42 97.34 $\pm$ 0.14 67.31 $\pm$ 0.31 0.435 $\pm$ 0.011
PNA - - - 97.19 $\pm$ 0.08 70.21 $\pm$ 0.15 0.320 $\pm$ 0.032
EGC-M - - - - 71.03 $\pm$ 0.42 0.281 $\pm$ 0.007
SIR-GCN 78.06 $\pm$ 0.66 85.75 $\pm$ 0.03 63.35 $\pm$ 0.19 97.90 $\pm$ 0.08 71.98 $\pm$ 0.40 0.278 $\pm$ 0.024
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al. [12] to ensure fairness. The results show that SIR-GCN consistently outperforms popular GNNs in literature. Notably, SIR-GCN also outperforms both PNA [11] and efficient graph convolution (EGC-M) [39] which use multiple aggregators. This highlights the significance of non-linear and contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
ogbn-arxiv
ogbn-arxiv [19] is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph [42]. Each node represents an arXiv paper and a directed edge represents a citation. The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
Table 4: Test accuracy on ogbn-arxiv.
Model GIANT-XRT [10] BoT [45] C&S [20] Others Accuracy Parameters
GATv2 0.7415 $\pm$ 0.0005 207,520
GraphSAGE 0.7435 $\pm$ 0.0014 546,344
SIR-GCN 0.7525 $\pm$ 0.0009 667,176
0.7574 $\pm$ 0.0020 697,896
LGGNN [30] 0.7570 $\pm$ 0.0018 1,161,640
DRGAT [1]
KD
0.7633 $\pm$ 0.0008 2,685,527
RevGAT [26]
KD, DCN
0.7636 $\pm$ 0.0013 1,304,912
AGDN [38]
self-KD
0.7637 $\pm$ 0.0011 1,309,760
Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers. Unsurprisingly, however, SIR-GCN performs poorly when compared against complex frameworks utilizing more tricks such as the dynamic evolving initial residual GAT (DRGAT) [1], reversible GAT (RevGAT) [26], and adaptive graph diffusion network (AGDN) [38], albeit with only a small performance difference.
ogbg-molhiv
ogbg-molhiv [19] is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds. Node features contain information regarding the atom while edge features contain information regarding the chemical bond. The task is to predict whether or not the molecules inhibit HIV replication.
Table 5: Test ROC-AUC on ogbg-molhiv.
Model GraphNorm [9] VirtualNode [14] Others ROC-AUC Parameters
GIN
FLAG
0.7748 $\pm$ 0.0096 3,336,306
GIN 0.7773 $\pm$ 0.0129 1,518,901
EGC-M 0.7818 $\pm$ 0.0153 317,265
GCN 0.7883 $\pm$ 0.0100 526,201
PNA 0.7905 $\pm$ 0.0132 326,081
SIR-GCN 0.7721 $\pm$ 0.0110 327,901
0.7981 $\pm$ 0.0062 328,201
GSN [6] 0.7799 $\pm$ 0.0100 3,338,701
GSAT [31] 0.8067 $\pm$ 0.0950 249,602
CIN [4] 0.8094 $\pm$ 0.0057 239,745
Table 5 presents the mean and standard deviation of the test ROC-AUC for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that with only a single GNN layer, SIR-GCN outperforms established models in predicting molecules inhibiting HIV replication, highlighting its expressivity. Nevertheless, SIR-GCN fails to compete with more complex models such as the graph stochastic attention (GSAT) [31] and models incorporating domain knowledge such as the cell isomorphism network (CIN) [4].
6 Conclusion
Overall, the paper provides a novel perspective for creating a powerful GNN when the space of node features is uncountable. The central idea is to use implicit pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This concept is demonstrated at all levels from node features to graph-level features. From the results, a novel SIR-GCN is proposed and shown to generalize classical GNN methodologies. The expressivity of SIR-GCN is then empirically demonstrated with synthetic and benchmark datasets, highlighting its relative superiority in outperforming comparable models. Future studies may explore incorporating SIR-GCN into existing frameworks such as using multiple aggregators [11], reversible GNNs [26], and adaptive graph diffusion networks [38] to achieve state-of-the-art performance.
Acknowledgments and Disclosure of Funding
The work is supported by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821 and 22K19833) and by Kyoto University and Toyota Motor Corporation through the joint project Advanced Mathematical Science for Mobility Society.
Appendix A Proofs
Definition 2 (Pseudometric).
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$.
  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
Definition 3 (Conditionally positive definite kernel [36]).
Let $\mathcal{H}$ be a non-empty set. A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$,
$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)\geq 0,$ (21)
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and
$\sum_{i=1}^{N}c_{i}=0.$ (22)
Lemma 1.
The negative of a pseudometric on $\mathcal{H}$ is a conditionally positive definite kernel on $\mathcal{H}$.
Proof.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. For all $N\in\mathbb{N}$ and $\bm{h}^{(0)},\bm{h}^{(1)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$, by the triangle inequality of $d$,
$-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)$ $\geq-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)$ (23)
$=-\sum_{j=1}^{N}c_{j}\sum_{i=1}^{N}c_{i}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}c_{i}\sum_{j=1}^{N}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)=0.$ (24)
Theorem 2 (Hilbert space representation of conditionally positive definite kernels [3, 35, 36]).
Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\bm{h},\bm{h}\right)=0$ for all $\bm{h}\in\mathcal{H}$. There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(1)}\in\mathcal{H}$,
$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (25)
Proof.
See Schölkopf [36]. ∎
See 1
Proof.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric. From Lemma 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
$d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}.$ (26)
For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}$ (27)
$\delta_{2}^{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}<\delta_{1}^{2}$ (28)
$\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (29)
Theorem 3.
Suppose $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
$k\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\dfrac{1}{2}\left[\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(0)}\right)-\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(2)}\right)+\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(0)}\right)\right]$ (30)
is positive definite if and only if $\tilde{k}$ is conditionally positive definite.
Proof.
See Schölkopf [36]. ∎
See 1
Proof.
Let $D$ be a pseudometric on bounded, equinumerous multisets of $\mathcal{H}$ defined as
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (31)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. From Lemma 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})+\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})-2\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}}).$ (32)
Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al. [23]. [^2] If $k$ is also integrally strictly positive definite [37], then the hash function $G$ becomes injective. By the linearity of the inner product, it may be shown that
$D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert,$ (33)
where
$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h})$ (34)
and $g$ is the corresponding feature map of the kernel $k$. For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
$\delta_{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert<\delta_{1}$ (35)
$\delta_{2}^{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert^{2}<\delta_{1}^{2}$ (36)
$\varepsilon_{2}<D\left(\bm{H_{1}},\bm{H_{2}}\right)<\varepsilon_{1}.$ (37)
Appendix B Experimental set-up
All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (dgl, version 2.1.0+cu121, Apache License 2.0) [43] with PyTorch (torch, version 2.2.1, BSD-3) [33] backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The hyperparameters are chosen based on previous results and heuristics without extensive tuning. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
B.1 Synthetic datasets
DictionaryLookup
Adopting Brody et al. [7], the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
GraphHeterophily
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN. Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
B.2 Benchmark datasets
Benchmarking GNNs
The dataset is obtained from dgl (version 2.1.0+cu121, Apache License 2.0) with data splits (training, validation, test) following Dwivedi et al. [12]. In line with Dwivedi et al. [12], all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. [41]. Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW [29] optimizer for over a maximum of 500 epochs with a batch size of 128 (whenever applicable) and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
ogbn-arxiv
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections. Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting. The symmetric mean aggregator is used along with existing tricks in literature. Model training is performed with the AdamW [29] optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
ogbg-molhiv
The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al. [19]. The models utilize a single GNN layer, modified to leverage edge features as described in Appendix C, with 300 hidden units, batch/graph normalization, and residual connections. Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting. The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function. Model training is performed with the AdamW [29] optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
Appendix C SIR-GCN extensions
Denote $\bm{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$. Following the intuition presented in Eqns. 8 and 9, SIR-GCN with residual connection may be modified to leverage edge features to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{E}}\bm{h_{u,v}}+\bm{W_{K}}\bm{h_{v}}\right),$ (38)
where $\bm{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (39)
with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features [18].
Furthermore, one may inject inductive bias into the pseudometric $D$ which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g. stock [17] and fMRI [24] data), $g$ may then be aptly modeled using recurrent-type networks.

### v3

@@
-Contextualized Messages Boost Graph Representations
+No License
 Abstract
 Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs.
-This success has prompted several studies to explore the representational capability of GNNs based on the graph isomorphism task.
+This has prompted several studies to explore their representational capability based on the graph isomorphism task.
 These works inherently assume a countable node feature representation, potentially limiting their applicability.
-Interestingly, only a few theoretical works study GNNs with uncountable node feature representation.
-This paper presents a novel perspective on the representational capability of GNNs across all levels – node-level, neighborhood-level, and graph-level – when the space of node feature representation is uncountable.
-Specifically, it relaxes the injective requirement in previous works by employing an implicit pseudometric distance on the space of input to create a soft-injective function.
-This allows distinct inputs to produce similar outputs only if the pseudometric deems the inputs to be sufficiently similar on some representation, which is often useful in practice.
-As a consequence, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes non-linear and contextualized transformation of neighborhood feature representations is proposed.
-A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution in context, establishing SIR-GCN as a generalization of classical GNN methodologies.
-Experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
+Interestingly, only a few study GNNs with uncountable node feature representation.
+In the paper, a novel perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable.
+More specifically, the strict injective and metric requirements are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation.
+As a consequence, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the contextualized transformation of neighborhood feature representations via anisotropic and dynamic message functions is proposed.
+A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies.
+Experiments on synthetic and benchmark datasets then demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
 1 Introduction
 Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs.
-These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others [12, 19].
-GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation [14, 47, 48].
+These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others (Dwivedi et al., 2023; Hu et al., 2020).
+GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018a; b).
 This allows the model to encapsulate both the network structure and the broader node contexts.
-Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph [27, 32, 47, 50].
-Among the most widely used GNNs in literature include the graph convolution network (GCN) [25], graph sample and aggregate (GraphSAGE) [15], graph attention network (GAT) [7, 40], and graph isomorphism network (GIN) [47] which largely fall under the message-passing neural networks (MPNNs) [14] framework.
-These models have gained popularity due to their simplicity and remarkable performance across various applications [12, 17, 19, 22, 24, 28].
-Improvements are also constantly being proposed to achieve state-of-the-art performance [4, 6, 21, 31, 38, 44, 49].
+Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph (Li et al., 2015; Murphy et al., 2019; Xu et al., 2018a; Ying et al., 2018).
+Among the most widely used GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2016), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Brody et al., 2021; Veličković et al., 2017), and graph isomorphism network (GIN) (Xu et al., 2018a) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework.
+These models have gained popularity due to their simplicity and remarkable performance across various applications (Dwivedi et al., 2023; Hsu et al., 2021; Hu et al., 2020; Jiang et al., 2022; Kim & Ye, 2020; Liu et al., 2020).
+Improvements are also constantly being proposed to achieve state-of-the-art performance (Bodnar et al., 2021; Bouritsas et al., 2022; Ishiguro et al., 2019; Miao et al., 2022; Sun et al., 2020; Wang et al., 2019b; Ying et al., 2021).
 Notably, these advances are mainly driven by heuristics and empirical results.
-Nonetheless, several studies have also begun exploring the representational capability of GNNs [2, 4, 5, 11, 13, 34].
+Nonetheless, several studies have also begun exploring the representational capability of GNNs (Azizian & Lelarge, 2020; Bodnar et al., 2021; Böker et al., 2024; Corso et al., 2020; Garg et al., 2020; Sato et al., 2021).
 Most of these works analyzed GNNs in relation to the graph isomorphism task.
 Xu et al.
-[47] was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test [46].
+(2018a) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968).
 Subsequent works build upon their results by considering extensions to the original 1-WL test.
@@
 Meanwhile, Corso et al.
-[11] proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable.
+(2020) proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable.
 Interestingly, there has been no significant theoretical progress since this work.
-This paper presents a simple yet novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable.
-The key idea is to define an implicit pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation.
-This idea is explored across all levels – node-level, neighborhood-level, and graph-level.
-Based on the results, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the non-linear and contextualized transformation of neighborhood feature representations is proposed.
-The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the advantages of the proposed model.
-Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN in closing.
+This paper presents a novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable.
+The key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation.
+This idea is investigated across all levels—node-level, neighborhood-level, and graph-level.
+From the theoretical results, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) which emphasizes the contextualized transformation of neighborhood feature representations using anisotropic and dynamic message functions is proposed.
+The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the novelty and advantages of the proposed model.
+Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN, positioning the proposed model as the best-performing MPNN instance.
 2 Graph neural networks
-Let $\mathcal{G}=\left(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{N}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{N}_{\mathcal{G}}$.
+Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$.
 The subscript $\mathcal{G}$ will be omitted whenever the context is clear.
@@
 $\bm{h^{*}_{u}}$ $=\textsc{COMB}\left(\bm{h_{u}},\bm{a_{u}}\right),$
-where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset [47] of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$.
+where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset (Xu et al., 2018a) of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$.
 Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function.
@@
 Related works
-When $\mathcal{H}$ is countable, Xu et al. [47] showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
+When $\mathcal{H}$ is countable, Xu et al. (2018a) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
 $F\left(\bm{H}\right)=\sum_{\bm{h}\in\bm{H}}f\left(\bm{h}\right)$ (2)
-is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size. This result forms the theoretical basis of GIN.
+is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size in the embedded feature space $\mathcal{S}$. This result forms the theoretical basis of GIN.
 Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable.
 In this setting, Corso et al.
-[11] proved that if $\bigoplus$ comprises multiple aggregators (e.g.
-mean, standard deviation, max, and min), the hash function
+(2020) proved that if $\bigoplus$ comprises multiple aggregators (e.g., mean, standard deviation, max, and min), the hash function
 $M\left(\bm{H}\right)=\bigoplus_{\bm{h}\in\bm{H}}m\left(\bm{h}\right)$ (3)
 produces a unique output for every $\bm{H}$ of bounded size.
-This finding laid the foundation for the principal neighborhood aggregation (PNA) [11].
+This finding provides the foundation for the principal neighborhood aggregation (PNA) (Corso et al., 2020).
 Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
 3 Soft-injective functions
-Theorem 1 presents an alternative to injective functions when the space of node features $\mathcal{H}$ is uncountable.
-It considers a soft relaxation to injectivity, henceforth soft-injectivity, which is often useful in practice, especially in classification tasks.
+While injective functions and metrics are necessary for tasks requiring strict isomorphism, many practical applications of GNNs often do not require such strict constraints.
+For instance, in node classification tasks, the model must produce identical outputs for some distinct nodes.
+Thus, this paper softly relaxes these constraints by employing pseudometrics and soft-injective functions.
+Definition 1 (Pseudometric).
+Let $\mathcal{H}$ be a non-empty set.
+A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$:
+  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
+  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
+  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
+[…]
+Assumption 1.
+Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric on $\mathcal{H}$ such that $-d^{2}$ is a conditionally positive definite kernel on $\mathcal{H}$.
+The Euclidean distance is an example of a pseudometric satisfying Assumption 1.
+A class of pseudometrics satisfying this assumption is provided below, see Berg et al.
+(1984) and Schölkopf (2000) for more.
+Remark 1.
+Consider the pseudometrics $d_{1}$ and $d_{2}$ on $\mathcal{H}$ satisfying Assumption 1.
+For $a>0$ and $10<p<$ 1, the pseudometrics $a\cdot d_{1}$, $\sqrt{d_{1}^{2}+d_{2}^{2}}$, and $d_{1}^{p}$ also satisfy Assumption 1.
+Assumption 1 thus offers considerable flexibility in the choice of pseudometric $d$. The following theorem then softly relaxes the injective and metric requirements in previous works.
 Theorem 1.
-Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$.
-There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
-$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
-Theorem 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function operating on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation.
-The lower bounds $\delta_{2}$ and $\varepsilon_{2}$ assert the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bounds $\delta_{1}$ and $\varepsilon_{1}$ ensure $g_{u}$ maintains the relationship of elements of $\mathcal{H}$ with respect to $d_{u}$.
-The feature map $g_{u}$ is then said to be soft-injective.
-An illustration is provided in Fig.
-1.
-$\delta$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
-(a) Embedded feature space $\mathcal{S}$.
+Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1.
+There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
 $d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)<\varepsilon<d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
-(b) Input feature space $\mathcal{H}$.
-Figure 1: Soft-injective feature map $g_{u}$ with pseudometric $d_{u}$.
-It is worth noting that if $d_{u}$ is squared Euclidean, then Theorem 1 becomes trivial.
-Nevertheless, the result becomes non-trivial for other choices of $d_{u}$.
+(a) Input feature space $\mathcal{H}$.
+$\varepsilon$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
+(b) Embedded feature space $\mathcal{S}$.
+Figure 1: Pseudometric $d_{u}$ and the corresponding feature map $g_{u}$.
+Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation.
+The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$.
+The feature map $g_{u}$ may then be described as soft-injective.
+[^1] The pseudometric $d$ induces the equivalence class $\left[\bm{h}\right]_{d}\coloneqq\left\{\bm{h^{\prime}}\in\mathcal{H}:d\left(\bm{h},\bm{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\bm{h}\right]_{d}:\bm{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938).
+Hence, $g$ may be described as soft-injective.
 Corollary 1 extends this result for multisets.
+3.1 Soft-isomorphic relational graph convolution network
 Corollary 1.
 Let $\mathcal{H}$ be a non-empty set with a pseudometric $D$ on bounded, equinumerous multisets of $\mathcal{H}$ defined as
-$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (5)
-for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$.
-There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
-$\delta_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
+$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (5)
+for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$.
+There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
 where
 $G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h}).$ (7)
-3.1 Soft-isomorphic relational graph convolution network
-Corollary 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn.
-5, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn.
-7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation.
-Similarly, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$.
-In this setting, the feature map $g_{u}$ may be interpreted as the message function [14] of the aggregation strategy that transforms the individual neighborhood features.
-Meanwhile, the distance function $D_{u}$ may be interpreted as a kernel distance [23].
-Definition 1 (Collision).
-Let $G$ be a function. If $G\left(\bm{H}^{(1)}\right)=G\left(\bm{H}^{(2)}\right)$ and $\bm{H}^{(1)}\neq\bm{H}^{(2)}$, a collision is said to have occurred.
-To illustrate the utility of $D_{u}$, suppose node $u$ has two neighbors $v_{1}$, $v_{2}$.
-If $d_{u}$ is the squared Euclidean distance, then a corresponding message function $g_{u}$ is linear.
-Fig.
-2(a) presents the contour plot of the corresponding hash function $G_{u}$, highlighting potential issues arising from hash collisions.
-Specifically, consider two multisets $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$.
-If $\bm{h_{v_{1}}}^{(1)}+\bm{h_{v_{2}}}^{(1)}=\bm{h_{v_{1}}}^{(2)}+\bm{h_{v_{2}}}^{(2)}$, then a hash collision occurs and $G_{u}$ produces identical aggregated neighborhood features even if $\bm{H_{u}}^{(1)}$ and $\bm{H_{u}}^{(2)}$ are fundamentally dissimilar on some representation for a given task.
+Similarly, Corollary 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn.
+5 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn.
+7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation.
+Likewise, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$.
+In this setting, the feature map $g_{u}$ may be interpreted as the message function (Gilmer et al., 2017) of the aggregation strategy that transforms the individual neighborhood features.
+Meanwhile, the psuedometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets.
+Dynamic transformation
+To illustrate the role of pseudometrics, consider anomaly detection on node features representing zero-mean scores, with node $u$ having two neighbors $v_{1}$ and $v_{2}$.
+If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as presented in Fig.
+2(a).
+The contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\bm{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
 […]
@@
 Figure 2: Hash functions $G_{u}$ under different message functions $g_{u}$.
-In general, hash collisions may occur when aggregating neighborhood features due to $\mathcal{H}$ being uncountable.
-Within the current framework, one may simply encode knowledge about node features into the pseudometric $D_{u}$.
-As a result, only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled.
-This also corresponds to a more complex and non-linear message function $g_{u}$.
-To illustrate, if node features represent a zero-mean score, $d_{u}$ may be defined as the squared Euclidean distance of the squared score.
-A corresponding hash function $G_{u}$ in Fig.
-2(b) may then be used to detect potentially anomalous neighborhoods since hash collisions are meaningful for this task.
-It is worth noting, however, that Corollary 1 holds for every node $u\in\mathcal{N}$ independently.
-Hence, different nodes may correspond to different $D_{u}$ and $G_{u}$.
-For simplicity, one may consider only a single pseudometric compactly defined as
-$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})$ (8)
-with a corresponding soft-injective hash function
-$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right)$ (9)
-for every node $u\in\mathcal{N}$.
-This approach preserves the interpretation of $G$ as an aggregation or hash function with an underlying pseudometric distance $D$ that guides and controls hash collisions.
-The integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function guiding how features of the key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$ are to be embedded and transformed based on the features of the query (center) node $\bm{h_{u}}$.
-This provides additional context to hash collisions and makes the message function and pseudometric anisotropic [12] or adaptive with respect to the query node.
+Nevertheless, other choices of pseudometrics, possibly incorporating prior knowledge, would correspond to more complex message functions $g_{u}$.
+This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled.
+This also highlights the significance of dynamic (Brody et al., 2021) or non-linear message functions $g_{u}$ in MPNNs.
+As further illustration, if $d_{u}$ instead corresponds to the Euclidean distance of the squared score, then the corresponding hash function $G_{u}$ has the contour plot in Fig.
+2(b).
+The resulting hash collisions and equivalence classes then become more useful and meaningful for detecting anomalous scores.
+Anisotropic messages
+It is also worth noting that Corollary 1 holds for each node $u\in\mathcal{V}$ independently.
+Hence, different nodes may correspond to different $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$.
+For simplicity, especially in inductive learning contexts, consider a single pseudometric instead, defined as
+$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}}),$ (8)
+with a single hash function, defined as
+$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right),$ (9)
+for every node $u\in\mathcal{V}$.
+This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\bm{h_{u}}$ and key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$).
+Moreover, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
+Furthermore, the integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node.
 Figs.
-2(c) and 2(d) demonstrate this idea where the introduction of a bias term, assuming a function of $\bm{h_{u}}$, shifts the contour plot and produces different aggregated neighborhood features for identical neighborhood features.
-Nevertheless, one may opt to inject stochasticity into the node features to distinguish between nodes with identical features and neighborhood features and imitate having distinct hash functions $G_{u}$ for every node $u$ with high probability [34].
-For a graph representation learning problem, the relational message function $g$ may be modeled as a two-layered multi-layer perceptron (MLP), with an implied pseudometric, following the universal approximation theorem [16] to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
+2(c) and 2(d) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\bm{a_{u}}\neq\bm{a_{u^{\prime}}}$ for nodes $u$ and $u^{\prime}$ with identical neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ but distinct features $\bm{h_{u}}\neq\bm{h_{u^{\prime}}}$.
+Nevertheless, one may also inject stochasticity into the node features to distinguish between nodes $u$ and $u^{\prime}$ with identical features $\bm{h_{u}}=\bm{h_{u^{\prime}}}$ and neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
+Proposed model
+For a graph representation learning problem, one may directly model the anisotropic and dynamic relational message function $g$ as a two-layer multi-layer perceptron (MLP), with implicitly learned pseudometrics, following the universal approximation theorem (Hornik et al., 1989) to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
 $\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (10)
@@
 Leveraging linearity, the model has a computational complexity of
-$\mathcal{O}\left(\left|\mathcal{N}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
-which is comparable to GAT.
-In practice, $\sigma$ may also be replaced with a deep MLP if modeling $g$ as a shallow two-layer MLP is infeasible.
-Moreover, since $\bm{h_{u}}$ is already encoded in the aggregation strategy, the combination strategy may simply be an activation function.
-In essence, the proposed SIR-GCN is an instance of the MPNN framework where, unlike most MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic [7] transformation of the neighborhood features to obtain contextualized messages.
+$\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (11)
+with computational efficiency achieved by the application of only an activation function along edges, making it comparable to GAT.
+Nevertheless, $\sigma$ may also be replaced with a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
+In essence, the proposed SIR-GCN is a simple, interpretable, and computationally efficient instance of the MPNN framework.
+Moreover, in contrast to other MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic transformation of neighborhood features to obtain contextualized messages.
 3.2 Soft-isomorphic graph readout function
-Corollary 1 also shows that, for every graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn.
+Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn.
 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn.
 7.
-While this result holds for every graph $\mathcal{G}$ independently, one may simply assume a single pseudometric $D$ with a corresponding soft-injective graph readout function $R$ for a set of graphs $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ from task $\mathcal{D}$.
-Nevertheless, one may opt to integrate additional information about the graph context and structure into the graph readout function $R$ to imitate having distinct $R_{\mathcal{G}}$ for every graph $\mathcal{G}$ and enhance its representational capability further.
-The virtual super node [14] may be used in this regard.
-In practice, $R$ may also be modeled with an MLP, with an implied pseudometric, to obtain the soft-isomorphic graph readout function
-$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{N}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
-where $\textsc{MLP}_{R}$ represents the corresponding feature map of $R$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
+While this result holds for each graph $\mathcal{G}$ independently, one may simply consider a single $D$, $d$, $r$, and $R$ for every graph $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ under task $\mathcal{D}$.
+Nevertheless, the graph context and structure may also be integrated into $D$, $d$, $r$, and $R$, through a virtual super node (Gilmer et al., 2017) for instance, to imitate having distinct $D_{\mathcal{G}}$, $d_{\mathcal{G}}$, $r_{\mathcal{G}}$, and $R_{\mathcal{G}}$ for each graph $\mathcal{G}$ and to further enhance its representational capability.
+Similarly, for a graph representation learning problem, $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
+$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
+where $\textsc{MLP}_{R}$ corresponds to $r$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
 4 Mathematical discussion
-The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the contribution.
-The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the proposed model.
+The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the novelty and contribution.
+While activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the message-passing operation that defines GNNs.
+The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
 4.1 GCN and GraphSAGE
 It may be shown that Corollary 1 holds up to a constant scale.
-Hence, the mean aggregation and symmetric mean aggregation (by extension) may be used in place of the sum aggregation.
+Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation.
 If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
@@
 [^2] GraphSAGE with LSTM aggregation is not included in this discussion.
-The difference lies in the isotropic [12] nature of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
+The difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
 4.2 GAT
-Moreover, in Brody et al. [7], the attention mechanism of GATv2 is modeled as an MLP given by
+Moreover, in Brody et al. (2021), the attention mechanism of GATv2 is modeled as an MLP given by
 $e_{u,v}=\bm{a}^{\top}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (15)
@@
 $\bm{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\bm{h_{u}}+\sum_{v\in\mathcal{N}(u)}\bm{h_{v}}$ (18)
-is equivalent to a GIN. Hence, SIR-GCN with residual connection encompasses GIN.
+is equivalent to GIN. Hence, SIR-GCN with residual connection encompasses GIN.
 4.4 PNA
-Furthermore, SIR-GCN and PNA approach the problem of uncountable node features $\mathcal{H}$ differently, with the former using only a single aggregator which holds theoretically for any number of neighbors.
-Nevertheless, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes.
-The key difference lies with PNA using a linear message function $m$ which is equivalent to a linear transformation of $\bm{h_{v}}$ with a different bias term for each $u$
+Furthermore, while SIR-GCN and PNA approach the problem of uncountable node features differently, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes.
+The key difference lies with PNA using a static (Brody et al., 2021) or linear message function $m$ which translates to
 $m\left(\bm{h_{v}},\bm{h_{u}}\right)=\bm{W_{K}}\bm{h_{v}}+\bm{W_{Q}}\bm{h_{u}}=\bm{W_{K}}\bm{h_{v}}+\bm{b_{u}}.$ (19)
-When using mean, max, or min aggregators, the query node $u$ only contributes to the aggregated neighborhood features through the bias term $\bm{b_{u}}$.
-Meanwhile, when using normalized moment aggregators, the query node $u$ no longer contributes to the aggregated neighborhood features.
-This potentially limits the expressivity of PNA since, in contrast to SIR-GCN and as suggested by Brody et al.
-[7], messages are only linearly and statically [7] transformed by the features of the query node.
+As a result, the influence of the query node on the aggregated neighborhood features is limited.
+For instance, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\bm{b_{u}}$.
+Moreover, with normalized moment aggregators, the bias term is effectively canceled out during the normalization process, further reducing the influence of the query node.
+Hence, PNA does not fully leverage its anisotropic nature, limiting its expressivity.
+In contrast, the dynamic nature of SIR-GCN allows for the non-linear embedding of the features of the query node $\bm{h_{u}}$ within the aggregated neighborhood features, thereby fully leveraging its anisotropic nature.
 4.5 1-WL test
@@
 Thus, SIR-GCN inherits the theoretical capabilities (and limitations) of the 1-WL test.
-Overall, SIR-GCN is demonstrated to generalize four prominent GNNs in literature – GCN, GraphSAGE, GAT, and GIN – and is thus at least as expressive as they are.
-Notably, SIR-GCN offers flexibility in two dimensions of GNNs: aggregation strategy and message transformation.
-It emphasizes summed aggregation with anisotropic and dynamic (i.e.
-contextualized) message transformation, making it well-suited for heterophilous tasks [8], but remains adaptable to alternative configurations.
-In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator, which theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity.
+4.6 SIR-GCN
+Overall, SIR-GCN offers flexibility in two key dimensions of GNNs: aggregation strategy and message transformation.
+Consequently, it generalizes four prominent GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models.
+Notably, SIR-GCN sets itself apart from other GNNs as the first MPNN instance to incorporate both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, making it well-suited for heterophilous tasks (Bronstein et al., 2021) while remaining adaptable to homophilous tasks.
+SumMeanAnisotropicIsotropicGATGINGraphSAGEGCNSIR-GCN
+Figure 3: SIR-GCN encompasses GCN, GraphSAGE, GAT, and GIN.
+In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity.
 Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
-SumMeanAnisotropicIsotropicGATGINGraphSAGEGCNSIR-GCN
-Figure 3: SIR-GCN expressivity.
 5 Experiments
 Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN.
-To ensure fairness, models not employing advanced architectural design or manually crafted features using domain knowledge are used as primary comparisons.
+To ensure fair evaluation, models not employing complex architectural design or manually crafted features using domain knowledge are used as primary comparisons.
 5.1 Synthetic datasets
 DictionaryLookup
-DictionaryLookup [7] consists of bipartite graphs with $2n$ nodes – $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute.
+DictionaryLookup (Brody et al., 2021) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute.
 The task is to predict the value of query nodes by matching their attribute with the key nodes.
-A sample instance is provided in Fig.
-4.
+$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
+Figure 4: DictionaryLookup.
+$A$$B$$A$$B$
+Figure 5: GraphHeterophily.
 Table 1: Test accuracy on DictionaryLookup.
@@
 GIN 0.78 $\pm$ 0.07 0.29 $\pm$ 0.03 0.12 $\pm$ 0.03 0.03 $\pm$ 0.00 0.02 $\pm$ 0.01
+PNA 1.00 $\pm$ 0.00 0.97 $\pm$ 0.02 0.86 $\pm$ 0.09 0.66 $\pm$ 0.09 0.50 $\pm$ 0.05
 SIR-GCN 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
-$A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
-Figure 4: DictionaryLookup.
-$A$$B$$A$$B$
-Figure 5: GraphHeterophily.
-Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $n$.
-SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic nature.
+Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $n$.
+SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic and dynamic nature.
 However, it is observed that GATv2 suffers from performance degradation in some trials.
-Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic nature.
-The results underscore the utility of an attentional/relational mechanism in capturing the relationship between the query and key nodes.
+Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature.
+The results underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
 GraphHeterophily
@@
 The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph.
-A sample instance is provided in Fig.
-5.
 Table 2: Test mean squared error on GraphHeterophily.
@@
 GIN 39.620 $\pm$ 2.060 37.193 $\pm$ 1.382 34.649 $\pm$ 1.502 32.424 $\pm$ 1.841 30.091 $\pm$ 1.429
+PNA 172.15 $\pm$ 97.82 224.83 $\pm$ 85.80 249.99 $\pm$ 108.56 251.49 $\pm$ 98.84 195.72 $\pm$ 36.65
 SIR-GCN 0.001 $\pm$ 0.000 0.004 $\pm$ 0.005 1.495 $\pm$ 4.428 0.038 $\pm$ 0.068 0.089 $\pm$ 0.134
-Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, and GIN across different values of $c$.
-SIR-GCN achieves an MSE loss of nearly 0 attributed to its anisotropic nature and sum aggregation.
+Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $c$.
+SIR-GCN achieves an MSE loss of nearly zero attributed to its anisotropic and dynamic nature and sum aggregation.
 In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al.
-[47].
-Meanwhile, GIN successfully retains the graph structure but fails to learn the relationship between the labels of the query (center) node and key (neighboring) nodes.
-The results illustrate the utility of anisotropic models even in graph property prediction tasks with countable node features.
+(2018a).
+Meanwhile, GIN and PNA successfully retain the graph structure but fail to learn the relationship between the labels of the query node and key nodes due to their static nature.
+The results illustrate the utility of anisotropic and dynamic models using sum aggregation even with countable node features.
 5.2 Benchmark datasets
 Benchmarking GNNs
-Benchmarking GNNs [12] is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks.
+Benchmarking GNNs (Dwivedi et al., 2023) is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks.
 In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks.
@@
 Dwivedi et al.
-[12] provides more information regarding the individual datasets.
+(2023) provides more information regarding the individual datasets.
 Table 3: Test performance on Benchmarking GNNs.
@@
 Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al.
-[12] to ensure fairness.
+(2023) to ensure fair evaluation.
 The results show that SIR-GCN consistently outperforms popular GNNs in literature.
-Notably, SIR-GCN also outperforms both PNA [11] and efficient graph convolution (EGC-M) [39] which use multiple aggregators.
-This highlights the significance of non-linear and contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
+Notably, SIR-GCN also outperforms both PNA (Corso et al., 2020) and efficient graph convolution (EGC-M) (Tailor et al., 2021) which use multiple aggregators.
+This highlights the significance of contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
 ogbn-arxiv
-ogbn-arxiv [19] is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph [42].
+ogbn-arxiv (Hu et al., 2020) is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph (Wang et al., 2020).
 Each node represents an arXiv paper and a directed edge represents a citation.
@@
 Table 4: Test accuracy on ogbn-arxiv.
-Model GIANT-XRT [10] BoT [45] C&S [20] Others Accuracy Parameters
+Model GIANT-XRT (Chien et al., 2021) BoT (Wang et al., 2021) C&S (Huang et al., 2020) Others Accuracy Parameters
 GATv2 0.7415 $\pm$ 0.0005 207,520
@@
 0.7574 $\pm$ 0.0020 697,896
-LGGNN [30] 0.7570 $\pm$ 0.0018 1,161,640
-DRGAT [1]
-KD
-0.7633 $\pm$ 0.0008 2,685,527
-RevGAT [26]
+LGGNN 0.7570 $\pm$ 0.0018 1,161,640
+RevGAT
 KD, DCN
 0.7636 $\pm$ 0.0013 1,304,912
-AGDN [38]
+AGDN
 self-KD
@@
 The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers.
-Unsurprisingly, however, SIR-GCN performs poorly when compared against complex frameworks utilizing more tricks such as the dynamic evolving initial residual GAT (DRGAT) [1], reversible GAT (RevGAT) [26], and adaptive graph diffusion network (AGDN) [38], albeit with only a small performance difference.
+As expected, however, SIR-GCN fails to compete with complex frameworks utilizing more tricks such as the reversible GAT (RevGAT) (Li et al., 2021) and the adaptive graph diffusion network (AGDN) (Sun et al., 2020), both of which build upon GAT by employing grouped reversible residual connections and adaptive graph diffusion, respectively.
+Nevertheless, SIR-GCN achieves performance close to that of the complex GNN frameworks mentioned, showcasing an effective balance between complexity and expressivity.
 ogbg-molhiv
-ogbg-molhiv [19] is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds.
+ogbg-molhiv (Hu et al., 2020) is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds.
 Node features contain information regarding the atom while edge features contain information regarding the chemical bond.
@@
 Table 5: Test ROC-AUC on ogbg-molhiv.
-Model GraphNorm [9] VirtualNode [14] Others ROC-AUC Parameters
+Model GraphNorm (Cai et al., 2021) VirtualNode (Gilmer et al., 2017) Others ROC-AUC Parameters
 GIN
@@
 0.7981 $\pm$ 0.0062 328,201
-GSN [6] 0.7799 $\pm$ 0.0100 3,338,701
-GSAT [31] 0.8067 $\pm$ 0.0950 249,602
-CIN [4] 0.8094 $\pm$ 0.0057 239,745
-Table 5 presents the mean and standard deviation of the test ROC-AUC for SIR-GCN and other models in literature.
+GSN 0.7799 $\pm$ 0.0100 3,338,701
+GSAT 0.8067 $\pm$ 0.0950 249,602
+CIN 0.8094 $\pm$ 0.0057 239,745
+Table 5 presents the mean and standard deviation of the test area under the receiver operating characteristic curve (ROC-AUC) for SIR-GCN and other models in literature.
 The tricks used and the number of parameters are also presented for completeness.
-The results show that with only a single GNN layer, SIR-GCN outperforms established models in predicting molecules inhibiting HIV replication, highlighting its expressivity.
-Nevertheless, SIR-GCN fails to compete with more complex models such as the graph stochastic attention (GSAT) [31] and models incorporating domain knowledge such as the cell isomorphism network (CIN) [4].
+The results show that with only a single GNN layer, SIR-GCN outperforms comparable models in predicting molecules inhibiting HIV replication, highlighting its expressivity.
+Given the simplicity of SIR-GCN, it is expected to exhibit lower performance compared to complex models such as the graph stochastic attention (GSAT) (Miao et al., 2022), which builds upon PNA by leveraging the information bottleneck principle, and the cell isomorphism network (CIN) (Bodnar et al., 2021), which is a hierarchical message-passing framework utilizing the topological features of graphs.
+Despite its simple design, the expressivity of SIR-GCN is evident in its close performance to that of complex GNN frameworks.
 6 Conclusion
-Overall, the paper provides a novel perspective for creating a powerful GNN when the space of node features is uncountable.
-The central idea is to use implicit pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation.
-This concept is demonstrated at all levels from node features to graph-level features.
-From the results, a novel SIR-GCN is proposed and shown to generalize classical GNN methodologies.
-The expressivity of SIR-GCN is then empirically demonstrated with synthetic and benchmark datasets, highlighting its relative superiority in outperforming comparable models.
-Future studies may explore incorporating SIR-GCN into existing frameworks such as using multiple aggregators [11], reversible GNNs [26], and adaptive graph diffusion networks [38] to achieve state-of-the-art performance.
-Acknowledgments and Disclosure of Funding
+In summary, the paper provides a novel perspective for creating a powerful GNN across all levels when the space of node features is uncountable.
+The central idea is to use pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between inputs is sufficiently small on some representation.
+From the results, the SIR-GCN is proposed as the first MPNN instance to emphasize contextualized message transformation, setting it apart from other GNNs.
+Furthermore, the model is shown to generalize classical GNN methodologies.
+Despite its simple design, empirical results underscore SIR-GCN as the best-performing MPNN instance that effectively balances model complexity and expressivity.
+The paper thus contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity.
+Future works may consider incorporating SIR-GCN into complex frameworks, such as grouped reversible residual connections (Li et al., 2021), adaptive graph diffusion (Sun et al., 2020), graph stochastic attention (Miao et al., 2022), and hierarchical message-passing (Bodnar et al., 2021), to address the limitations inherent in the MPNN framework and to develop more expressive GNN architectures.
+Acknowledgments
 The work is supported by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821 and 22K19833) and by Kyoto University and Toyota Motor Corporation through the joint project Advanced Mathematical Science for Mobility Society.
 Appendix A Proofs
-Definition 2 (Pseudometric).
-Let $\mathcal{H}$ be a non-empty set.
-A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$.
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
-Definition 3 (Conditionally positive definite kernel [36]).
+Definition 2 (Conditionally positive definite kernel (Schölkopf, 2000)).
 Let $\mathcal{H}$ be a non-empty set.
@@
 $\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)\geq 0,$ (21)
-with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and
-$\sum_{i=1}^{N}c_{i}=0.$ (22)
-Lemma 1.
-The negative of a pseudometric on $\mathcal{H}$ is a conditionally positive definite kernel on $\mathcal{H}$.
-Proof.
-Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric.
-For all $N\in\mathbb{N}$ and $\bm{h}^{(0)},\bm{h}^{(1)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$, by the triangle inequality of $d$,
-$-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)$ $\geq-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)$ (23)
-$=-\sum_{j=1}^{N}c_{j}\sum_{i=1}^{N}c_{i}~{}d\left(\bm{h}^{(i)},\bm{h}^{(0)}\right)-\sum_{i=1}^{N}c_{i}\sum_{j=1}^{N}c_{j}~{}d\left(\bm{h}^{(0)},\bm{h}^{(j)}\right)=0.$ (24)
-Theorem 2 (Hilbert space representation of conditionally positive definite kernels [3, 35, 36]).
+with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
+Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Berg et al., 1984; Schoenberg, 1938; Schölkopf, 2000)).
 Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\bm{h},\bm{h}\right)=0$ for all $\bm{h}\in\mathcal{H}$.
@@
 Proof.
-See Schölkopf [36]. ∎
+See Schölkopf (2000). ∎
 See 1
 Proof.
-Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ be a pseudometric.
-From Lemma 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
-$d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}.$ (26)
-For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
-$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}$ (27)
-$\delta_{2}^{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}<\delta_{1}^{2}$ (28)
-$\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (29)
+Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric.
+From Assumption 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
+$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert=d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (23)
+Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
 Theorem 3.
@@
 Proof.
-See Schölkopf [36]. ∎
+See Schölkopf (2000). ∎
 See 1
@@
 Let $D$ be a pseudometric on bounded, equinumerous multisets of $\mathcal{H}$ defined as
-$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (31)
-for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$.
-From Lemma 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$.
+$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (26)
+for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$.
+From Assumption 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$.
 A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
@@
 Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al.
-[23].
-[^2] If $k$ is also integrally strictly positive definite [37], then the hash function $G$ becomes injective.
-By the linearity of the inner product, it may be shown that
-$D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert,$ (33)
+(2011).
+[^3] If $k$ is also integrally strictly positive definite (Sriperumbudur et al., 2010), then the hash function $G$ becomes injective and $D$ becomes a metric.
+By the reproducing property of $k$ and the linearity of the inner product, it may be shown that
+$\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert=D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right),$ (28)
 where
 $G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h})$ (29)
-and $g$ is the corresponding feature map of the kernel $k$. For $\varepsilon_{1}>\varepsilon_{2}>0$, let $\delta_{1}=\sqrt{\varepsilon_{1}},\delta_{2}=\sqrt{\varepsilon_{2}}$. Hence,
-$\delta_{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert<\delta_{1}$ (35)
-$\delta_{2}^{2}<\left\lVert G\left(\bm{H_{1}}\right)-G\left(\bm{H_{2}}\right)\right\rVert^{2}<\delta_{1}^{2}$ (36)
-$\varepsilon_{2}<D\left(\bm{H_{1}},\bm{H_{2}}\right)<\varepsilon_{1}.$ (37)
+and $g$ is the corresponding feature map of the kernel $k$. Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
 Appendix B Experimental set-up
-All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (dgl, version 2.1.0+cu121, Apache License 2.0) [43] with PyTorch (torch, version 2.2.1, BSD-3) [33] backend.
+All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend.
 For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values.
@@
 Adopting Brody et al.
-[7], the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration.
+(2021), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration.
 All models utilize a single GNN layer with $4n$ hidden units.
-A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN.
-Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and std aggregators.
+Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
 GraphHeterophily
@@
 All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function.
-A feed-forward neural network is also used for GIN.
-Model training is performed with the AdamW [29] optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+A feed-forward neural network is also used for GIN while PNA uses the sum, max, and std aggregators.
+Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
 B.2 Benchmark datasets
 Benchmarking GNNs
-The dataset is obtained from dgl (version 2.1.0+cu121, Apache License 2.0) with data splits (training, validation, test) following Dwivedi et al.
-[12].
+The datasets are obtained from dgl with data splits (training, validation, test) following Dwivedi et al.
+(2023).
 In line with Dwivedi et al.
-[12], all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000.
+(2023), all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000.
 Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting.
 The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al.
-[41].
+(2019).
 Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10.
-Model training is performed with the AdamW [29] optimizer for over a maximum of 500 epochs with a batch size of 128 (whenever applicable) and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+The reported results for other models in Table 3 are obtained from Dwivedi et al.
+(2023), Corso et al.
+(2020), and Tailor et al.
+(2021).
 ogbn-arxiv
-The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al.
-[19].
+The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al.
+(2020).
 The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections.
@@
 The symmetric mean aggregator is used along with existing tricks in literature.
-Model training is performed with the AdamW [29] optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
+Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
+The reported results for other models in Table 4 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
 ogbg-molhiv
-The dataset is obtained from ogb (version 1.3.6, MIT License) with data splits (training, validation, test) following Hu et al.
-[19].
-The models utilize a single GNN layer, modified to leverage edge features as described in Appendix C, with 300 hidden units, batch/graph normalization, and residual connections.
+The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al.
+(2020).
+The models utilize a single GNN layer, modified to leverage edge features as described in Appendix D, with 300 hidden units, batch/graph normalization, and residual connections.
 Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting.
 The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function.
-Model training is performed with the AdamW [29] optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
-Appendix C SIR-GCN extensions
+Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
+The reported results for other models in Table 5 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
+Appendix C Model run time
+As an additional evaluation, the run time per epoch for each model in the synthetic datasets is presented in Tables 6 and 7.
+The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity.
+Notably, SIR-GCN is computationally less expensive than PNA while delivering superior performance.
+These results further underscore the practicality of the proposed model.
+Table 6: Run time per epoch on DictionaryLookup.
+Model $n=10$ $n=20$ $n=30$ $n=40$ $n=50$
+GCN 0.3526s $\pm$ 0.0778s 0.4734s $\pm$ 0.0468s 0.4777s $\pm$ 0.0854s 0.5619s $\pm$ 0.0518s 0.5520s $\pm$ 0.0679s
+GraphSAGE 0.4565s $\pm$ 0.0873s 0.5264s $\pm$ 0.0317s 0.5716s $\pm$ 0.1132s 0.7742s $\pm$ 0.0597s 0.9193s $\pm$ 0.0473s
+GATv2 0.3950s $\pm$ 0.1017s 0.5276s $\pm$ 0.0556s 0.6191s $\pm$ 0.0879s 0.7472s $\pm$ 0.0346s 1.0065s $\pm$ 0.0280s
+GIN 0.3696s $\pm$ 0.0899s 0.4610s $\pm$ 0.0459s 0.4670s $\pm$ 0.0781s 0.5947s $\pm$ 0.0548s 0.5194s $\pm$ 0.0993s
+PNA 0.8854s $\pm$ 0.0412s 1.1913s $\pm$ 0.1024s 1.4526s $\pm$ 0.0684s 1.8793s $\pm$ 0.0528s 2.8387s $\pm$ 0.0603s
+SIR-GCN 0.4687s $\pm$ 0.0777s 0.6066s $\pm$ 0.0398s 0.8053s $\pm$ 0.0485s 1.1496s $\pm$ 0.0427s 1.7031s $\pm$ 0.0458s
+Table 7: Run time per epoch on GraphHeterophily.
+Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
+GCN 0.4243s $\pm$ 0.0520s 0.3852s $\pm$ 0.0517s 0.3868s $\pm$ 0.0743s 0.4166s $\pm$ 0.0551s 0.4177s $\pm$ 0.0494s
+GraphSAGE 0.4691s $\pm$ 0.0400s 0.4790s $\pm$ 0.0440s 0.4399s $\pm$ 0.0629s 0.4501s $\pm$ 0.0603s 0.4964s $\pm$ 0.0601s
+GATv2 0.4710s $\pm$ 0.0978s 0.4941s $\pm$ 0.0567s 0.4718s $\pm$ 0.0361s 0.5514s $\pm$ 0.0608s 0.5437s $\pm$ 0.0724s
+GIN 0.4085s $\pm$ 0.0741s 0.3875s $\pm$ 0.0627s 0.3855s $\pm$ 0.0645s 0.4298s $\pm$ 0.0566s 0.4329s $\pm$ 0.0534s
+PNA 2.2963s $\pm$ 0.0413s 2.4238s $\pm$ 0.0611s 2.4577s $\pm$ 0.0533s 2.4741s $\pm$ 0.0665s 2.5623s $\pm$ 0.0425s
+SIR-GCN 0.5338s $\pm$ 0.0353s 0.5264s $\pm$ 0.0737s 0.5635s $\pm$ 0.0695s 0.5764s $\pm$ 0.0401s 0.6230s $\pm$ 0.0388s
+Appendix D SIR-GCN extensions
 Denote $\bm{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$.
@@
 where $\bm{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
-$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{N}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{N}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (39)
+$\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (32)
 with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA.
-Similarly, this extension may be viewed as a generalization of GIN with edge features [18].
-Furthermore, one may inject inductive bias into the pseudometric $D$ which may correspond to specifying the architecture type for the corresponding message function $g$.
-For instance, if node features are known to have a sequential relationship (e.g.
-stock [17] and fMRI [24] data), $g$ may then be aptly modeled using recurrent-type networks.
+Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2019).
+Furthermore, one may inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$.
+For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2021) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.

### v4

@@
-No License
+Contextualized Messages Boost Graph Representations
 Abstract
@@
 This has prompted several studies to explore their representational capability based on the graph isomorphism task.
-These works inherently assume a countable node feature representation, potentially limiting their applicability.
+Notably, these works inherently assume a countable node feature representation, potentially limiting their applicability.
 Interestingly, only a few study GNNs with uncountable node feature representation.
-In the paper, a novel perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable.
-More specifically, the strict injective and metric requirements are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation.
+In the paper, a new perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable.
+Specifically, the injective and metric requirements of previous works are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation.
 As a consequence, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the contextualized transformation of neighborhood feature representations via anisotropic and dynamic message functions is proposed.
-A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies.
-Experiments on synthetic and benchmark datasets then demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
+Furthermore, a mathematical discussion on the relationship between SIR-GCN and key GNNs in literature is laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies.
+To close, experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
 1 Introduction
 Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs.
-These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others (Dwivedi et al., 2023; Hu et al., 2020).
-GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018a; b).
+These models are well-suited for node, edge, and graph property prediction tasks across various domains, including social networks, molecular graphs, and biological networks, among others (Hu et al., 2020a; Dwivedi et al., 2023).
+In the literature, GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018; 2019).
 This allows the model to encapsulate both the network structure and the broader node contexts.
-Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph (Li et al., 2015; Murphy et al., 2019; Xu et al., 2018a; Ying et al., 2018).
-Among the most widely used GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2016), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Brody et al., 2021; Veličković et al., 2017), and graph isomorphism network (GIN) (Xu et al., 2018a) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework.
-These models have gained popularity due to their simplicity and remarkable performance across various applications (Dwivedi et al., 2023; Hsu et al., 2021; Hu et al., 2020; Jiang et al., 2022; Kim & Ye, 2020; Liu et al., 2020).
-Improvements are also constantly being proposed to achieve state-of-the-art performance (Bodnar et al., 2021; Bouritsas et al., 2022; Ishiguro et al., 2019; Miao et al., 2022; Sun et al., 2020; Wang et al., 2019b; Ying et al., 2021).
-Notably, these advances are mainly driven by heuristics and empirical results.
-Nonetheless, several studies have also begun exploring the representational capability of GNNs (Azizian & Lelarge, 2020; Bodnar et al., 2021; Böker et al., 2024; Corso et al., 2020; Garg et al., 2020; Sato et al., 2021).
+Moreover, a graph readout function is employed to pool the node feature representations of a graph and create an aggregated representation for the entire graph (Ying et al., 2018; Murphy et al., 2019; Xu et al., 2019).
+Among the classical GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2017), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Veličković et al., 2018; Brody et al., 2022), and graph isomorphism network (GIN) (Xu et al., 2019) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework.
+These models have gained popularity due to their simplicity and remarkable performance across various applications (Hu et al., 2020a; Dwivedi et al., 2023).
+Improvements of these foundational models are also constantly proposed to achieve state-of-the-art performance (Wang et al., 2019b; Bodnar et al., 2021; Bouritsas et al., 2023).
+Notably, advances in GNN have mainly been driven by heuristics and empirical results.
+Nonetheless, several studies have recently begun exploring the representational capability of GNNs (Garg et al., 2020; Sato et al., 2021; Azizian & Lelarge, 2021; Bodnar et al., 2021; Böker et al., 2023).
 Most of these works analyzed GNNs in relation to the graph isomorphism task.
-Xu et al.
-(2018a) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968).
+In particular, Xu et al.
+(2019) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968).
 Subsequent works build upon their results by considering extensions to the original 1-WL test.
-However, these results only hold with countable node feature representation which potentially limits their applicability.
+Crucially, the theoretical results of these works only hold with countable node feature representation which potentially limits their applicability.
 Meanwhile, Corso et al.
@@
 Interestingly, there has been no significant theoretical progress since this work.
-This paper presents a novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable.
-The key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation.
-This idea is investigated across all levels—node-level, neighborhood-level, and graph-level.
+This paper presents a new perspective on the representational capability of GNNs when the space of node feature representation is uncountable.
+Specifically, the key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation.
+This framework is comprehensively analyzed across all levels—node-level, neighborhood-level, and graph-level.
 From the theoretical results, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) which emphasizes the contextualized transformation of neighborhood feature representations using anisotropic and dynamic message functions is proposed.
-The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the novelty and advantages of the proposed model.
-Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN, positioning the proposed model as the best-performing MPNN instance.
+This is further accompanied by a discussion on the mathematical relationship between SIR-GCN and key GNNs in literature to underscore the contribution and distinctive advantages of the proposed model.
+Finally, experiments on synthetic and benchmark datasets in node and graph property prediction tasks are performed to highlight the expressivity of SIR-GCN, positioning the model as a strong candidate for practical GNN applications.
 2 Graph neural networks
-Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$.
+Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ be the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$.
 The subscript $\mathcal{G}$ will be omitted whenever the context is clear.
-Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$.
+Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\boldsymbol{h_{u}}\in\mathcal{H}$ is the feature of node $u$.
 A GNN following the message-passing scheme can be expressed mathematically as
-$\bm{H_{u}}$ =$=\left\{\!\!\left\{\bm{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}$
-$\bm{a_{u}}$ $=\textsc{AGG}\left(\bm{H_{u}}\right)$ (1)
-$\bm{h^{*}_{u}}$ $=\textsc{COMB}\left(\bm{h_{u}},\bm{a_{u}}\right),$
-where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset (Xu et al., 2018a) of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$.
+$\boldsymbol{H_{u}}$ :-$\coloneq\left\{\!\!\left\{\boldsymbol{h_{v}}:v\in\mathcal{N}_{\mathcal{G}}(u)\right\}\!\!\right\}:$
+$\boldsymbol{a_{u}}$ :-$\coloneq\textsc{AGG}\left(\boldsymbol{H_{u}}\right):$ (1)
+$\boldsymbol{h^{*}_{u}}$ :-$\coloneq\textsc{COMB}\left(\boldsymbol{h_{u}},\boldsymbol{a_{u}}\right),:$
+where AGG and COMB are some aggregation and combination strategies, respectively, $\boldsymbol{H_{u}}$ is the multiset (Xu et al., 2019) of neighborhood features for node $u$, $\boldsymbol{a_{u}}$ is the aggregated neighborhood feature for node $u$, and $\boldsymbol{h^{*}_{u}}$ is the updated feature for node $u$.
 Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function.
@@
 Related works
-When $\mathcal{H}$ is countable, Xu et al. (2018a) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
-$F\left(\bm{H}\right)=\sum_{\bm{h}\in\bm{H}}f\left(\bm{h}\right)$ (2)
-is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size in the embedded feature space $\mathcal{S}$. This result forms the theoretical basis of GIN.
+When $\mathcal{H}$ is countable, Xu et al. (2019) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
+$F\left(\boldsymbol{H}\right)\coloneq\sum_{\boldsymbol{h}\in\boldsymbol{H}}f\left(\boldsymbol{h}\right)$ (2)
+is injective or unique in the embedded feature space $\mathcal{S}$ for each multiset of neighborhood features $\boldsymbol{H}$ of bounded size. This result forms the theoretical basis of GIN.
 Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable.
@@
 (2020) proved that if $\bigoplus$ comprises multiple aggregators (e.g., mean, standard deviation, max, and min), the hash function
-$M\left(\bm{H}\right)=\bigoplus_{\bm{h}\in\bm{H}}m\left(\bm{h}\right)$ (3)
-produces a unique output for every $\bm{H}$ of bounded size.
+$M\left(\boldsymbol{H}\right)\coloneq\bigoplus_{\boldsymbol{h}\in\boldsymbol{H}}m\left(\boldsymbol{h}\right)$ (3)
+produces a unique output for every $\boldsymbol{H}$ of bounded size.
 This finding provides the foundation for the principal neighborhood aggregation (PNA) (Corso et al., 2020).
-Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
+Notably, for this result to hold theoretically, the number of aggregators in $\bigoplus$ must scale with the size of the multiset of neighborhood features $\boldsymbol{H}$ which may be impractical for large and dense graphs.
 3 Soft-injective functions
-While injective functions and metrics are necessary for tasks requiring strict isomorphism, many practical applications of GNNs often do not require such strict constraints.
-For instance, in node classification tasks, the model must produce identical outputs for some distinct nodes.
-Thus, this paper softly relaxes these constraints by employing pseudometrics and soft-injective functions.
+While injective functions and metrics are necessary for tasks requiring graph isomorphism to ensure a unique mapping in the embedded feature space, many practical applications of GNN often do not require such strict constraints.
+For instance, in node classification tasks, GNNs must produce identical outputs for some distinct nodes.
+Thus, motivated by Xu et al.
+(2019) and Corso et al.
+(2020), this paper softly relaxes the injective and metric requirements within the MPNN framework by employing pseudometrics and soft-injective functions.
 Definition 1 (Pseudometric).
 Let $\mathcal{H}$ be a non-empty set.
-A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$:
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=d\left(\bm{h}^{(2)},\bm{h}^{(1)}\right)$; and
-  • $d\left(\bm{h}^{(1)},\bm{h}^{(3)}\right)\leq d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)+d\left(\bm{h}^{(2)},\bm{h}^{(3)}\right)$.
-[…]
+A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)},\boldsymbol{h}^{(3)}\in\mathcal{H}$:
+  1. 1.
+    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(1)}\right)=0$;
+  2. 2.
+    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=d\left(\boldsymbol{h}^{(2)},\boldsymbol{h}^{(1)}\right)$; and
+  3. 3.
+    $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(3)}\right)\leq d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)+d\left(\boldsymbol{h}^{(2)},\boldsymbol{h}^{(3)}\right)$.
+Note that for a metric $d$, the first condition is replaced with $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=0\Longleftrightarrow\boldsymbol{h}^{(1)}=\boldsymbol{h}^{(2)}$, ensuring points in $\mathcal{H}$ are distinguishable and unique with respect to $d$.
+In contrast, for a pseudometric $d$, $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=0$ does not necessarily imply that $\boldsymbol{h}^{(1)}=\boldsymbol{h}^{(2)}$, relaxing the distinguishability constraint of a metric.
+The following assumption is then imposed on the pseudometric $d$, leveraging results from kernel theory.
+Definition 2 (Conditionally positive definite kernel (Schölkopf, 2000)).
+Let $\mathcal{H}$ be a non-empty set.
+A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)},\ldots,\boldsymbol{h}^{(N)}\in\mathcal{H}$,
+$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\boldsymbol{h}^{(i)},\boldsymbol{h}^{(j)}\right)\geq 0,$ (4)
+with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
 Assumption 1.
@@
 The Euclidean distance is an example of a pseudometric satisfying Assumption 1.
-A class of pseudometrics satisfying this assumption is provided below, see Berg et al.
-(1984) and Schölkopf (2000) for more.
+A class of pseudometrics satisfying this assumption is provided below, see Schölkopf (2000) and Berg et al.
+(2012) for more.
 Remark 1.
@@
 Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1.
-There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
-$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
-$d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)<\varepsilon<d_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}},{\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
-(a) Input feature space $\mathcal{H}$.
-$\varepsilon$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
-(b) Embedded feature space $\mathcal{S}$.
-Figure 1: Pseudometric $d_{u}$ and the corresponding feature map $g_{u}$.
-Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation.
-The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$.
+There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)<\varepsilon_{1}.$ (5)
+Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\boldsymbol{h_{u}}^{(1)},\boldsymbol{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\boldsymbol{h_{u}}^{(1)}$ and $\boldsymbol{h_{u}}^{(2)}$ to be sufficiently similar on some representation.
+The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures that $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$.
 The feature map $g_{u}$ may then be described as soft-injective.
-[^1] The pseudometric $d$ induces the equivalence class $\left[\bm{h}\right]_{d}\coloneqq\left\{\bm{h^{\prime}}\in\mathcal{H}:d\left(\bm{h},\bm{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\bm{h}\right]_{d}:\bm{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938).
-Hence, $g$ may be described as soft-injective.
+[^1] The pseudometric $d$ induces the equivalence class $\left[\boldsymbol{h}\right]_{d}\coloneqq\left\{\boldsymbol{h^{\prime}}\in\mathcal{H}:d\left(\boldsymbol{h},\boldsymbol{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\boldsymbol{h}\right]_{d}:\boldsymbol{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938).
+Fig.
+1 provides an illustration of how the soft-injective feature map $g$ maps distinct elements of $\mathcal{H}$ to the same point in $\mathcal{S}$ since the corresponding pseudometric $d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)=\left\lVert\left[\boldsymbol{h}^{(1)}\right]^{2}-\left[\boldsymbol{h}^{(2)}\right]^{2}\right\rVert$ deems these points to be similar.
 Corollary 1 extends this result for multisets.
+$\left(\mathcal{H},d\right)$${\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\boldsymbol{h^{(1)}}}$${\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\boldsymbol{h^{(2)}}}$${\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\boldsymbol{h^{(3)}}}$$\left(\mathcal{H},d\right)$$\mathcal{S}$$g(d)$${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\mathcal{H}}$$\mathcal{S}$$g\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\boldsymbol{h^{(1)}}}\right)$$g\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\boldsymbol{h^{(2)}}}\right)$$g\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\boldsymbol{h^{(3)}}}\right)$
+Figure 1: A soft-injective feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ corresponding to a pseudometric $d$ on $\mathcal{H}$.
 3.1 Soft-isomorphic relational graph convolution network
@@
 Let $\mathcal{H}$ be a non-empty set with a pseudometric $D$ on bounded, equinumerous multisets of $\mathcal{H}$ defined as
-$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (5)
-for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$.
-There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
-$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
+$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})$ (6)
+for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$.
+There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
+$\varepsilon_{2}<\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)<\varepsilon_{1},$ (7)
 where
-$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h}).$ (7)
+$G(\boldsymbol{H})=\sum_{\boldsymbol{h}\in\boldsymbol{H}}g(\boldsymbol{h}).$ (8)
 Similarly, Corollary 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn.
-5 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn.
-7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation.
+6 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and hash function $G_{u}$ defined in Eqn.
+8 that produces similar outputs for distinct multisets of neighborhood features $\boldsymbol{H_{u}}^{(1)},\boldsymbol{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\boldsymbol{H_{u}}^{(1)}$ and $\boldsymbol{H_{u}}^{(2)}$ to be sufficiently similar on some representation.
 Likewise, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$.
-In this setting, the feature map $g_{u}$ may be interpreted as the message function (Gilmer et al., 2017) of the aggregation strategy that transforms the individual neighborhood features.
-Meanwhile, the psuedometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets.
+In this setting, the feature map $g_{u}$ may be interpreted as the soft-injective message function (Gilmer et al., 2017) that transforms the individual neighborhood features with a corresponding soft-injective hash function $G_{u}$.
+Meanwhile, the pseudometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets.
+The two necessary properties of the soft-injective message function—dynamic and anisotropic—are then motivated below.
 Dynamic transformation
-To illustrate the role of pseudometrics, consider anomaly detection on node features representing zero-mean scores, with node $u$ having two neighbors $v_{1}$ and $v_{2}$.
-If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as presented in Fig.
+To illustrate the role of the pseudometric, consider node $u$ with two neighbors $v_{1}$ and $v_{2}$ and the task of anomaly detection on the scalar node features $\boldsymbol{h_{v_{1}}}$ and $\boldsymbol{h_{v_{2}}}$ representing zero-mean scores.
+If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as seen in Fig.
 2(a).
-The contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\bm{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
+Crucially, the contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\boldsymbol{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
 […]
-(a) $g_{u}(\bm{h})=\bm{h}$.
+(a) $g_{u}(\boldsymbol{h})=\boldsymbol{h}$.
 […]
-(b) $g_{u}(\bm{h})=-\bm{h}^{2}$.
+(b) $g_{u}(\boldsymbol{h})=\textsc{MLP}(\boldsymbol{h})$.
 […]
-(c) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h})$.
+(c) $g_{u}(\boldsymbol{h})=\textsc{MLP}(\boldsymbol{h}+\boldsymbol{1})$.
 […]
-(d) $g_{u}(\bm{h})=\textsc{MLP}(\bm{h}+\bm{1})$.
+(d) $g_{u}(\boldsymbol{h})=-\boldsymbol{h}^{2}$.
 Figure 2: Hash functions $G_{u}$ under different message functions $g_{u}$.
 Nevertheless, other choices of pseudometrics, possibly incorporating prior knowledge, would correspond to more complex message functions $g_{u}$.
-This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled.
-This also highlights the significance of dynamic (Brody et al., 2021) or non-linear message functions $g_{u}$ in MPNNs.
+This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar with respect to a given task may produce similar aggregated neighborhood features, making collisions more informative and controlled.
+In practice, this also highlights the significance of dynamic (Brody et al., 2022) (i.e., a universal function approximator (Hornik et al., 1989)) message functions $g_{u}$ in the MPNN framework, which may be modeled as multi-layer perceptrons (MLPs) as illustrated in Figs.
+2(b) and 2(c).
 As further illustration, if $d_{u}$ instead corresponds to the Euclidean distance of the squared score, then the corresponding hash function $G_{u}$ has the contour plot in Fig.
-2(b).
+2(d).
 The resulting hash collisions and equivalence classes then become more useful and meaningful for detecting anomalous scores.
@@
 For simplicity, especially in inductive learning contexts, consider a single pseudometric instead, defined as
-$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}}),$ (8)
+$D^{2}\left(\boldsymbol{H_{u}}^{(1)},\boldsymbol{H_{u}}^{(2)};\boldsymbol{h_{u}}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}}),$ (9)
 with a single hash function, defined as
-$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right),$ (9)
+$G\left(\boldsymbol{H_{u}};\boldsymbol{h_{u}}\right)\coloneq\sum_{\boldsymbol{h}\in\boldsymbol{H_{u}}}g\left(\boldsymbol{h};\boldsymbol{h_{u}}\right),$ (10)
 for every node $u\in\mathcal{V}$.
-This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\bm{h_{u}}$ and key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$).
-Moreover, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
-Furthermore, the integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node.
+This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\boldsymbol{h_{u}}$ and key (neighboring) nodes $\boldsymbol{h}\in\boldsymbol{H_{u}}$).
+In addition, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
+Furthermore, the integration of $\boldsymbol{h_{u}}$ also allows for the interpretation of $g$ as a soft-injective relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node.
 Figs.
-2(c) and 2(d) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\bm{a_{u}}\neq\bm{a_{u^{\prime}}}$ for nodes $u$ and $u^{\prime}$ with identical neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ but distinct features $\bm{h_{u}}\neq\bm{h_{u^{\prime}}}$.
-Nevertheless, one may also inject stochasticity into the node features to distinguish between nodes $u$ and $u^{\prime}$ with identical features $\bm{h_{u}}=\bm{h_{u^{\prime}}}$ and neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
+2(b) and 2(c) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\boldsymbol{a_{u}}\neq\boldsymbol{a_{u^{\prime}}}$ for nodes $u\neq u^{\prime}\in\mathcal{V}$ with identical neighborhood features $\boldsymbol{H_{u}}=\boldsymbol{H_{u^{\prime}}}$ but distinct features $\boldsymbol{h_{u}}\neq\boldsymbol{h_{u^{\prime}}}$.
+Moreover, one may also inject stochasticity into the node features to distinguish between nodes $u\neq u^{\prime}\in\mathcal{V}$ with identical features $\boldsymbol{h_{u}}=\boldsymbol{h_{u^{\prime}}}$ and neighborhood features $\boldsymbol{H_{u}}=\boldsymbol{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
 Proposed model
-For a graph representation learning problem, one may directly model the anisotropic and dynamic relational message function $g$ as a two-layer multi-layer perceptron (MLP), with implicitly learned pseudometrics, following the universal approximation theorem (Hornik et al., 1989) to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
-$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (10)
-where $\sigma$ is a non-linear activation function, $\bm{W_{Q}},\bm{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\bm{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$.
+For a graph representation learning problem, one may directly model the anisotropic and dynamic soft-injective relational message function $g$ as a two-layer MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
+$\boldsymbol{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (11)
+where $\sigma$ is a non-linear activation function, $\boldsymbol{W_{Q}},\boldsymbol{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\boldsymbol{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$.
 Leveraging linearity, the model has a computational complexity of
 $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ (12)
-with computational efficiency achieved by the application of only an activation function along edges, making it comparable to GAT.
-Nevertheless, $\sigma$ may also be replaced with a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
-In essence, the proposed SIR-GCN is a simple, interpretable, and computationally efficient instance of the MPNN framework.
-Moreover, in contrast to other MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic transformation of neighborhood features to obtain contextualized messages.
+with only the application of an activation function along edges, making it comparable to classical GNNs in literature.
+Nevertheless, $\sigma$ may also be modeled as a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
+In essence, SIR-GCN is a simple instance of the MPNN framework explicitly designed to handle uncountable node features while maintaining rigorous theoretical foundations.
+It emphasizes the anisotropic and dynamic transformation of neighborhood features, obtaining contextualized messages that enable it to learn complex relationships between neighboring nodes.
+Moreover, the proposed model is also computationally efficient, requiring only a single aggregator and applying only an activation function along edges to facilitate effective message-passing of uncountable node features.
 3.2 Soft-isomorphic graph readout function
-Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn.
-5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn.
-7.
+Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn.
+6 with a corresponding pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$, there exists a corresponding feature map $r_{\mathcal{G}}$ and graph readout function $R_{\mathcal{G}}$ defined in Eqn.
+8.
 While this result holds for each graph $\mathcal{G}$ independently, one may simply consider a single $D$, $d$, $r$, and $R$ for every graph $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ under task $\mathcal{D}$.
 Nevertheless, the graph context and structure may also be integrated into $D$, $d$, $r$, and $R$, through a virtual super node (Gilmer et al., 2017) for instance, to imitate having distinct $D_{\mathcal{G}}$, $d_{\mathcal{G}}$, $r_{\mathcal{G}}$, and $R_{\mathcal{G}}$ for each graph $\mathcal{G}$ and to further enhance its representational capability.
-Similarly, for a graph representation learning problem, $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
-$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
-where $\textsc{MLP}_{R}$ corresponds to $r$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
+Similarly, for a graph representation learning problem, the dynamic soft-injective feature map $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
+$\boldsymbol{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\boldsymbol{h_{v}}\right),$ (13)
+where $\textsc{MLP}_{R}$ corresponds to $r$ and $\boldsymbol{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
 4 Mathematical discussion
-The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the novelty and contribution.
-While activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the message-passing operation that defines GNNs.
-The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
+The mathematical relationship between SIR-GCN and key GNNs in literature—GCN, GraphSAGE, GAT, GIN, and PNA—are presented in this section to underscore the contribution and distinctive advantages of the proposed model.
+It is worth noting that while activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the core message-passing operation that defines GNNs.
+In addition, the relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
 4.1 GCN and GraphSAGE
 It may be shown that Corollary 1 holds up to a constant scale.
-Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation.
-If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
-$\bm{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\bm{W}\bm{h_{v}}$ (13)
+Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation in Eqn.
+11.
+If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\boldsymbol{W_{Q}}=\boldsymbol{0}$, $\boldsymbol{W_{R}}\boldsymbol{W_{K}}=\boldsymbol{W}$, and $\tilde{\mathcal{N}}(u)\coloneq\mathcal{N}(u)\cup\{u\}$, one obtains
+$\boldsymbol{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\boldsymbol{W}\boldsymbol{h_{v}}$ (14)
 and
-$\bm{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\bm{W}\bm{h_{v}}$ (14)
-which recovers GCN and GraphSAGE with mean aggregation, respectively.
+$\boldsymbol{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\boldsymbol{W}\boldsymbol{h_{v}}$ (15)
+which are equivalent to GCN and GraphSAGE with mean aggregation, respectively.
 Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling.
-Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN.
-[^2] GraphSAGE with LSTM aggregation is not included in this discussion.
-The difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.
+Thus, GCN and GraphSAGE [^2] GraphSAGE with LSTM aggregation is not included in this discussion.
+may be viewed as instances of SIR-GCN.
+The key difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and their non-linearities only after aggregating the neighborhood features.
 4.2 GAT
 Moreover, in Brody et al. (2022), the attention mechanism of GATv2 is modeled as an MLP given by
-$e_{u,v}=\bm{a}^{\top}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (15)
-with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\bm{W}\bm{h_{v}}$.
-While the model is anisotropic in nature, messages are nevertheless only linearly transformed with node $u$ only determining the degree of contribution through the scalar $e_{u,v}$.
-Meanwhile, SIR-GCN directly works with the unnormalized attention mechanism in Eqn.
-15 and allows the features of the query node to dynamically transform messages.
-Specifically, if $\sigma=\textsc{LeakyReLU}$ and $\bm{W_{R}}=\bm{A}$, one obtains
-$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{A}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right)$ (16)
-which shows Eqn. 15 becoming a contextualized message in the SIR-GCN model. Nevertheless, GAT and GATv2 may be recovered, up to a normalizing constant, with an appropriate choice of $\sigma$.
+$e_{u,v}=\boldsymbol{a_{\text{GAT}}^{\top}}~{}\textsc{LeakyReLU}\left(\boldsymbol{W_{Q,\text{GAT}}}~{}\boldsymbol{h_{u}}+\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}\right),$ (16)
+with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}$.
+While the attention mechanism of GATv2 is anisotropic and dynamic, its messages are nevertheless only linearly transformed with the query node $u$ only determining the degree of contribution of each message through the scalar $e_{u,v}$.
+Meanwhile, SIR-GCN directly applies the anisotropic and dynamic function in Eqn.
+16 to the message function, allowing the features of the query node to dynamically transform messages.
+Specifically, if $\boldsymbol{W_{Q}}=\boldsymbol{W_{Q,\text{GAT}}}$, $\boldsymbol{W_{K}}=\boldsymbol{W_{K,\text{GAT}}}$, $\sigma=\textsc{LeakyReLU}$ and $\boldsymbol{W_{R}}=\boldsymbol{a_{\text{GAT}}^{\top}}$, one obtains
+$\boldsymbol{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\boldsymbol{a_{\text{GAT}}^{\top}}~{}\textsc{LeakyReLU}\left(\boldsymbol{W_{Q,\text{GAT}}}~{}\boldsymbol{h_{u}}+\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}\right)$ (17)
+which shows Eqn. 16 becoming a contextualized message in SIR-GCN. Nevertheless, GAT and GATv2 may also be recovered, up to a normalizing constant, with the appropriate parameters.
 4.3 GIN
-Likewise, within the proposed SIR-GCN model, one may explicitly add a residual connection in the combination strategy to obtain
-$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (17)
+Likewise, within the proposed SIR-GCN, one may explicitly add a residual connection in the combination strategy to obtain
+$\boldsymbol{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\boldsymbol{h_{u}})+\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (18)
 where $\textsc{MLP}_{\text{Res}}$ is a learnable residual network.
-If $\textsc{MLP}_{\text{Res}}(\bm{h})=(1+\epsilon)\cdot\bm{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, and $\bm{W_{R}W_{K}}=\bm{I}$, then
-$\bm{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\bm{h_{u}}+\sum_{v\in\mathcal{N}(u)}\bm{h_{v}}$ (18)
-is equivalent to GIN. Hence, SIR-GCN with residual connection encompasses GIN.
+If $\textsc{MLP}_{\text{Res}}(\boldsymbol{h})=(1+\epsilon)\cdot\boldsymbol{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\boldsymbol{W_{Q}}=\boldsymbol{0}$, and $\boldsymbol{W_{R}W_{K}}=\boldsymbol{I}$, then
+$\boldsymbol{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\boldsymbol{h_{u}}+\sum_{v\in\mathcal{N}(u)}\boldsymbol{h_{v}}$ (19)
+is equivalent to GIN. Hence, SIR-GCN with residual connection generalizes GIN.
 4.4 PNA
 Furthermore, while SIR-GCN and PNA approach the problem of uncountable node features differently, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes.
-The key difference lies with PNA using a static (Brody et al., 2021) or linear message function $m$ which translates to
-$m\left(\bm{h_{v}},\bm{h_{u}}\right)=\bm{W_{K}}\bm{h_{v}}+\bm{W_{Q}}\bm{h_{u}}=\bm{W_{K}}\bm{h_{v}}+\bm{b_{u}}.$ (19)
-As a result, the influence of the query node on the aggregated neighborhood features is limited.
-For instance, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\bm{b_{u}}$.
+The key difference lies with PNA using multiple aggregators and a static (Brody et al., 2022) (i.e., a function approximator with limited expressivity; e.g., linear) message function
+$m\left(\boldsymbol{h_{v}},\boldsymbol{h_{u}}\right)=\boldsymbol{W_{K}}\boldsymbol{h_{v}}+\boldsymbol{W_{Q}}\boldsymbol{h_{u}}\eqcolon\boldsymbol{W_{K}}\boldsymbol{h_{v}}+\boldsymbol{b_{u}}.$ (20)
+Consequently, the influence of the query node on the aggregated neighborhood feature becomes limited.
+In particular, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\boldsymbol{b_{u}}\coloneq\boldsymbol{W_{Q}}\boldsymbol{h_{u}}$.
 Moreover, with normalized moment aggregators, the bias term is effectively canceled out during the normalization process, further reducing the influence of the query node.
-Hence, PNA does not fully leverage its anisotropic nature, limiting its expressivity.
-In contrast, the dynamic nature of SIR-GCN allows for the non-linear embedding of the features of the query node $\bm{h_{u}}$ within the aggregated neighborhood features, thereby fully leveraging its anisotropic nature.
+Hence, PNA does not fully leverage its anisotropic nature, attributed to its heuristic application of multiple aggregators and scalers in a linear MPNN, thereby limiting its expressivity.
+In contrast, the dynamic nature of SIR-GCN allows for the non-linear and contextualized embedding of the features of the query node within the messages, thereby fully leveraging its anisotropic nature while allowing it to employ only a single aggregator.
 4.5 1-WL test
@@
 $w^{(l)}_{u}\leftarrow\text{hash}\left(\left\{\!\!\!\left\{\left[w^{(l-1)}_{v},w^{(l-1)}_{u}\right]:v\in\mathcal{N}(u)\right\}\!\!\!\right\}\right),$ (21)
-where the modification lies in concatenating the label of the center node with every element of the multiset before hashing.
-This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as noted in the previous section.
-Thus, SIR-GCN inherits the theoretical capabilities (and limitations) of the 1-WL test.
+where the modification lies in concatenating the label of the query node $u$ with every element of the multiset before hashing.
+This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as highlighted in the previous section.
+Thus, SIR-GCN inherits the theoretical capabilities of the 1-WL test.
 4.6 SIR-GCN
-Overall, SIR-GCN offers flexibility in two key dimensions of GNNs: aggregation strategy and message transformation.
-Consequently, it generalizes four prominent GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models.
-Notably, SIR-GCN sets itself apart from other GNNs as the first MPNN instance to incorporate both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, making it well-suited for heterophilous tasks (Bronstein et al., 2021) while remaining adaptable to homophilous tasks.
+Overall, SIR-GCN is a simple MPNN instance that offers flexibility in two key dimensions of GNN—message transformation and aggregator.
+Consequently, it generalizes four classical GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models.
+Notably, SIR-GCN distinguishes itself from other GNNs by employing both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, enabling the non-uniform aggregation of neighboring nodes in heterophilous graphs (Zheng et al., 2024) while maintaining adaptability to homophilous graphs.
 SumMeanAnisotropicIsotropicGATGINGraphSAGEGCNSIR-GCN
-Figure 3: SIR-GCN encompasses GCN, GraphSAGE, GAT, and GIN.
-In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity.
-Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
+Figure 3: SIR-GCN generalizes classical GNNs.
+In addition, SIR-GCN also distinguishes itself from PNA in addressing the problem of uncountable node features by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity.
+Nevertheless, its expressivity is maintained through contextualized messages via the application of only an activation function along edges, allowing it to inherit the representational capability of the 1-WL test.
 5 Experiments
-Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN.
-To ensure fair evaluation, models not employing complex architectural design or manually crafted features using domain knowledge are used as primary comparisons.
+Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are performed to highlight the expressivity of SIR-GCN.
+Following the evaluation methodology of Xu et al.
+(2019), Corso et al.
+(2020), and Brody et al.
+(2022), the key GNNs in literature without advanced architectural design nor domain-specific features are used as primary comparisons to ensure a fair evaluation.
 5.1 Synthetic datasets
 DictionaryLookup
-DictionaryLookup (Brody et al., 2021) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute.
-The task is to predict the value of query nodes by matching their attribute with the key nodes.
+DictionaryLookup (Brody et al., 2022) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute.
+The task is to predict the values of query nodes by matching their attributes with the key nodes as seen in Fig.
+4.
 $A,1$$B,2$$C,3$$D,4$$A,*$$B,*$$C,*$$D,*$
@@
 $A$$B$$A$$B$
-Figure 5: GraphHeterophily.
+Figure 5: HeteroEdgeCount.
 Table 1: Test accuracy on DictionaryLookup.
@@
 Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $n$.
-SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic and dynamic nature.
-However, it is observed that GATv2 suffers from performance degradation in some trials.
-Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature.
-The results underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
-GraphHeterophily
-GraphHeterophily is an original synthetic dataset.
-It consists of random directed graphs with each node assigned one of $c$ classes.
-The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph.
-Table 2: Test mean squared error on GraphHeterophily.
+Notably, SIR-GCN and GATv2 can achieve perfect accuracy in this synthetic task attributed to their anisotropic and dynamic nature, enabling them to learn the relationship between every query and key node.
+Nonetheless, it may be observed that GATv2 suffers from performance degradation in some trials.
+Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature, hindering their ability to learn relationships between neighboring nodes.
+The results hence underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
+HeteroEdgeCount
+HeteroEdgeCount is an original synthetic dataset consisting of randomly generated directed graphs with each node randomly labeled one of $c$ classes.
+The task is then to count the number of heterophilous directed edges in each graph connecting nodes with different class labels as illustrated in Fig.
+5.
+This dataset is explicitly designed to highlight the limitations of key GNNs in literature, particularly the theoretically grounded GNNs such as GIN and PNA, even in trivial tasks involving countable node features.
+Specifically, it underscores the utility of anisotropic and dynamic message functions in learning the relationships between neighboring nodes.
+Crucially, this dataset is not intended to assess the ability of GNNs to handle heterophilous graphs, as this falls beyond the scope of this work.
+Table 2: Test mean squared error on HeteroEdgeCount.
 Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
@@
 Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $c$.
-SIR-GCN achieves an MSE loss of nearly zero attributed to its anisotropic and dynamic nature and sum aggregation.
+Notably, SIR-GCN consistently achieves near-zero MSE loss due to its anisotropic and dynamic nature as well as its sum aggregation, allowing it to learn the relationship between the labels of neighboring nodes while retaining the graph structure.
+In fact, for $\boldsymbol{W_{Q}}=\boldsymbol{I}$, $\boldsymbol{W_{K}}=-\boldsymbol{I}$, $\sigma=\textsc{ReLU}$, and $\boldsymbol{W_{R}}=\boldsymbol{1}^{\top}$, it may be shown that SIR-GCN will always produce the correct output for every graph.
 In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al.
-(2018a).
-Meanwhile, GIN and PNA successfully retain the graph structure but fail to learn the relationship between the labels of the query node and key nodes due to their static nature.
-The results illustrate the utility of anisotropic and dynamic models using sum aggregation even with countable node features.
+(2019).
+Meanwhile, GIN and PNA successfully retain the graph structure with their sum aggregation but fail to differentiate neighboring nodes due to their static nature.
+The results thus illustrate the utility of anisotropic and dynamic message functions using sum aggregation even in simple tasks with countable node features, highlighting the limitations of existing GNNs.
 5.2 Benchmark datasets
@@
 Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features.
-The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy.
+The performance metric of ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy.
+These six benchmark datasets encompass a diverse range of GNN tasks, enabling a comprehensive and robust evaluation of model performance.
 Dwivedi et al.
@@
 Model WikiCS ($\uparrow$) PATTERN ($\uparrow$) CLUSTER ($\uparrow$) MNIST ($\uparrow$) CIFAR10 ($\uparrow$) ZINC ($\downarrow$)
-MLP 59.45 $\pm$ 2.33 50.52 $\pm$ 0.00 20.97 $\pm$ 0.00 95.34 $\pm$ 0.14 56.34 $\pm$ 0.18 0.706 $\pm$ 0.006
 GCN 77.47 $\pm$ 0.85 85.50 $\pm$ 0.05 47.83 $\pm$ 1.51 90.12 $\pm$ 0.15 54.14 $\pm$ 0.39 0.416 $\pm$ 0.006
@@
 GAT 76.91 $\pm$ 0.82 75.82 $\pm$ 1.82 57.73 $\pm$ 0.32 95.54 $\pm$ 0.21 64.22 $\pm$ 0.46 0.475 $\pm$ 0.007
+GATv2 - - - - 67.48 $\pm$ 0.53 0.447 $\pm$ 0.015
 GIN 75.86 $\pm$ 0.58 85.59 $\pm$ 0.01 58.38 $\pm$ 0.24 96.49 $\pm$ 0.25 55.26 $\pm$ 1.53 0.387 $\pm$ 0.015
-GatedGCN - 84.48 $\pm$ 0.12 60.40 $\pm$ 0.42 97.34 $\pm$ 0.14 67.31 $\pm$ 0.31 0.435 $\pm$ 0.011
 PNA - - - 97.19 $\pm$ 0.08 70.21 $\pm$ 0.15 0.320 $\pm$ 0.032
+EGC-S - - - - 66.92 $\pm$ 0.37 0.364 $\pm$ 0.020
 EGC-M - - - - 71.03 $\pm$ 0.42 0.281 $\pm$ 0.007
 SIR-GCN 78.06 $\pm$ 0.66 85.75 $\pm$ 0.03 63.35 $\pm$ 0.19 97.90 $\pm$ 0.08 71.98 $\pm$ 0.40 0.278 $\pm$ 0.024
-Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al.
-(2023) to ensure fair evaluation.
-The results show that SIR-GCN consistently outperforms popular GNNs in literature.
-Notably, SIR-GCN also outperforms both PNA (Corso et al., 2020) and efficient graph convolution (EGC-M) (Tailor et al., 2021) which use multiple aggregators.
-This highlights the significance of contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
-ogbn-arxiv
-ogbn-arxiv (Hu et al., 2020) is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph (Wang et al., 2020).
-Each node represents an arXiv paper and a directed edge represents a citation.
-The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
-Table 4: Test accuracy on ogbn-arxiv.
-Model GIANT-XRT (Chien et al., 2021) BoT (Wang et al., 2021) C&S (Huang et al., 2020) Others Accuracy Parameters
-GATv2 0.7415 $\pm$ 0.0005 207,520
-GraphSAGE 0.7435 $\pm$ 0.0014 546,344
-SIR-GCN 0.7525 $\pm$ 0.0009 667,176
-0.7574 $\pm$ 0.0020 697,896
-LGGNN 0.7570 $\pm$ 0.0018 1,161,640
-RevGAT
-KD, DCN
-0.7636 $\pm$ 0.0013 1,304,912
-AGDN
-self-KD
-0.7637 $\pm$ 0.0011 1,309,760
-Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature.
-The tricks used and the number of parameters are also presented for completeness.
-The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers.
-As expected, however, SIR-GCN fails to compete with complex frameworks utilizing more tricks such as the reversible GAT (RevGAT) (Li et al., 2021) and the adaptive graph diffusion network (AGDN) (Sun et al., 2020), both of which build upon GAT by employing grouped reversible residual connections and adaptive graph diffusion, respectively.
-Nevertheless, SIR-GCN achieves performance close to that of the complex GNN frameworks mentioned, showcasing an effective balance between complexity and expressivity.
-ogbg-molhiv
-ogbg-molhiv (Hu et al., 2020) is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds.
-Node features contain information regarding the atom while edge features contain information regarding the chemical bond.
-The task is to predict whether or not the molecules inhibit HIV replication.
-Table 5: Test ROC-AUC on ogbg-molhiv.
-Model GraphNorm (Cai et al., 2021) VirtualNode (Gilmer et al., 2017) Others ROC-AUC Parameters
-GIN
-FLAG
-0.7748 $\pm$ 0.0096 3,336,306
-GIN 0.7773 $\pm$ 0.0129 1,518,901
-EGC-M 0.7818 $\pm$ 0.0153 317,265
-GCN 0.7883 $\pm$ 0.0100 526,201
-PNA 0.7905 $\pm$ 0.0132 326,081
-SIR-GCN 0.7721 $\pm$ 0.0110 327,901
-0.7981 $\pm$ 0.0062 328,201
-GSN 0.7799 $\pm$ 0.0100 3,338,701
-GSAT 0.8067 $\pm$ 0.0950 249,602
-CIN 0.8094 $\pm$ 0.0057 239,745
-Table 5 presents the mean and standard deviation of the test area under the receiver operating characteristic curve (ROC-AUC) for SIR-GCN and other models in literature.
-The tricks used and the number of parameters are also presented for completeness.
-The results show that with only a single GNN layer, SIR-GCN outperforms comparable models in predicting molecules inhibiting HIV replication, highlighting its expressivity.
-Given the simplicity of SIR-GCN, it is expected to exhibit lower performance compared to complex models such as the graph stochastic attention (GSAT) (Miao et al., 2022), which builds upon PNA by leveraging the information bottleneck principle, and the cell isomorphism network (CIN) (Bodnar et al., 2021), which is a hierarchical message-passing framework utilizing the topological features of graphs.
-Despite its simple design, the expressivity of SIR-GCN is evident in its close performance to that of complex GNN frameworks.
+Note: Missing values indicate that no results were published in previous works.
+Table 3 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, GAT, GATv2, GIN, and PNA across the six benchmark datasets with the experimental set-up, such as parameter count and model architecture, following that of Dwivedi et al.
+(2023), Corso et al.
+(2020), and Tailor et al.
+(2022) to ensure a fair evaluation where performance differences are solely attributed to the GNN architectural design.
+The test performance for the efficient graph convolution single (EGC-S) and efficient graph convolution multiple (EGC-M) (Tailor et al., 2022) are also presented as additional MPNN-based baselines.
+Notably, SIR-GCN consistently outperforms classical GNNs in literature by a substantial margin which may be attributed to its ability to generalize these models, complementing the mathematical discussions in the previous section.
+Moreover, despite employing multiple aggregators and incurring higher computational complexity to ensure injectivity, PNA still fails to outperform the simpler and more computationally efficient SIR-GCN on datasets with uncountable node features, even though the former is explicitly designed for such tasks.
+Furthermore, SIR-GCN also outperforms the more recent EGC-S and EGC-M despite their use of additional tricks, including multiple convolutional basis weights, regularization heads, and aggregators.
+Overall, the results underscore that, under the same constraints, SIR-GCN consistently outperforms MPNN-based baselines despite its simplicity, establishing it as a promising alternative to existing MPNNs.
+Open Graph Benchmark
+Open Graph Benchmark (Hu et al., 2020a) is another collection of datasets consisting of realistic, large-scale, and diverse benchmarks for GNNs.
+In particular, the ogbn-arxiv with its uncountable node features falls under node property prediction tasks.
+Meanwhile, the ogbg-molhiv with its countable node features falls under graph property prediction tasks.
+The performance metric of ogbn-arxiv is accuracy while the performance metric of ogbg-molhiv is the area under the receiver operating characteristic curve (ROC-AUC).
+Hu et al.
+(2020a) provides more information regarding the individual datasets.
+Table 4: Test performance on Open Graph Benchmark.
+Model ogbn-arxiv ($\uparrow$) ogbg-molhiv ($\uparrow$)
+GCN 71.92 $\pm$ 0.21 76.14 $\pm$ 1.29
+GraphSAGE 71.73 $\pm$ 0.26 75.97 $\pm$ 1.69
+GAT 71.81 $\pm$ 0.23 77.17 $\pm$ 1.37
+GATv2 71.87 $\pm$ 0.43 77.15 $\pm$ 1.55
+GIN 67.33 $\pm$ 1.47 76.02 $\pm$ 1.35
+PNA 71.21 $\pm$ 0.30 79.05 $\pm$ 1.32
+EGC-S 72.21 $\pm$ 0.17 77.44 $\pm$ 1.08
+EGC-M 71.96 $\pm$ 0.23 78.18 $\pm$ 1.53
+SIR-GCN 72.52 $\pm$ 0.16 77.63 $\pm$ 0.84
+Table 4 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, GAT, GATv2, GIN, PNA, EGC-S, and EGC-M across the two large-scale benchmark datasets with the experimental set-up following that of Corso et al.
+(2020) and Tailor et al.
+(2022) to ensure a fair evaluation.
+Notably, SIR-GCN still outperforms both classical GNNs and EGC-S by a significant margin even in these large-scale graphs, which may be attributed to its anisotropic and dynamic message function.
+Unsurprisingly, however, PNA and EGC-M exhibit better performance relative to SIR-GCN on ogbg-molhiv as this molecular property prediction task greatly benefits from maintaining graph isomorphism via multiple aggregators, scalers, convolutional basis weights, and regularization heads.
+Nevertheless, SIR-GCN outperforms both PNA and EGC-M by a substantial margin on ogbn-arxiv as this 40-class node classification task does not require maintaining graph isomorphism.
+Overall, the results highlight the significance of contextualized messages in enhancing GNN expressivity and the utility of softly relaxing the injective and metric requirements within the MPNN framework for most practical GNN applications.
 6 Conclusion
-In summary, the paper provides a novel perspective for creating a powerful GNN across all levels when the space of node features is uncountable.
-The central idea is to use pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between inputs is sufficiently small on some representation.
-From the results, the SIR-GCN is proposed as the first MPNN instance to emphasize contextualized message transformation, setting it apart from other GNNs.
-Furthermore, the model is shown to generalize classical GNN methodologies.
-Despite its simple design, empirical results underscore SIR-GCN as the best-performing MPNN instance that effectively balances model complexity and expressivity.
-The paper thus contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity.
-Future works may consider incorporating SIR-GCN into complex frameworks, such as grouped reversible residual connections (Li et al., 2021), adaptive graph diffusion (Sun et al., 2020), graph stochastic attention (Miao et al., 2022), and hierarchical message-passing (Bodnar et al., 2021), to address the limitations inherent in the MPNN framework and to develop more expressive GNN architectures.
+In summary, the paper provides a new perspective for creating powerful GNNs when the space of node features is uncountable.
+The key idea is to use pseudometric distances on the space of input to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation.
+From the theoretical results, SIR-GCN is proposed as a simple and computationally efficient MPNN instance emphasizing contextualized message transformation.
+Notably, compared to existing MPNN instances, this distinctive feature enables it to learn complex relationships between neighboring nodes and allows it to better handle uncountable node features.
+Furthermore, the proposed model is also demonstrated to generalize classical GNN methodologies.
+Despite its simple architectural design and minimal computational requirements, empirical results on synthetic and benchmark datasets underscore the expressivity of SIR-GCN, making it a promising candidate for practical GNN applications.
+Overall, the paper contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity.
+Future works may extend the present framework by considering more complex pseudometric formulations for bounded, equinumerous multisets of $\mathcal{H}$ in Corollary 1.
+They may also consider a formal analysis of the relationship between contextualized messages and performance on heterophilous graph tasks.
 Acknowledgments
-The work is supported by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821 and 22K19833) and by Kyoto University and Toyota Motor Corporation through the joint project Advanced Mathematical Science for Mobility Society.
+This work is supported in part by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821) and in part by Kyoto University and Toyota Motor Corporation through the joint project titled “Advanced Mathematical Science for Mobility Society.”
 Appendix A Proofs
-Definition 2 (Conditionally positive definite kernel (Schölkopf, 2000)).
-Let $\mathcal{H}$ be a non-empty set.
-A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$,
-$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)\geq 0,$ (21)
-with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
-Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Berg et al., 1984; Schoenberg, 1938; Schölkopf, 2000)).
-Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\bm{h},\bm{h}\right)=0$ for all $\bm{h}\in\mathcal{H}$.
-There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(1)}\in\mathcal{H}$,
-$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (22)
+Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Schoenberg, 1938; Schölkopf, 2000; Berg et al., 2012)).
+Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\boldsymbol{h},\boldsymbol{h}\right)=0$ for all $\boldsymbol{h}\in\mathcal{H}$.
+There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(1)}\in\mathcal{H}$,
+$\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right).$ (22)
 Proof.
@@
 Proof.
-Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric.
-From Assumption 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
-$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert=d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (23)
+Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric satisfying Assumption 1.
+By Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$,
+$\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert=d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right).$ (23)
 Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
-$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
+$\varepsilon_{2}<\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
 Theorem 3.
-Suppose $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
-$k\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\dfrac{1}{2}\left[\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(0)}\right)-\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(2)}\right)+\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(0)}\right)\right]$ (25)
+Suppose $\boldsymbol{h}^{(0)},\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
+$k\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)\coloneq\dfrac{1}{2}\left[\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)-\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(0)}\right)-\tilde{k}\left(\boldsymbol{h}^{(0)},\boldsymbol{h}^{(2)}\right)+\tilde{k}\left(\boldsymbol{h}^{(0)},\boldsymbol{h}^{(0)}\right)\right]$ (25)
 is positive definite if and only if $\tilde{k}$ is conditionally positive definite.
@@
 Let $D$ be a pseudometric on bounded, equinumerous multisets of $\mathcal{H}$ defined as
-$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (26)
-for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$.
-From Assumption 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$.
-A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
-$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})+\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})-2\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}}).$ (27)
+$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})$ (26)
+for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$.
+By Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$.
+A simple algebraic manipulation and using the fact that $\boldsymbol{H}^{(1)}$ and $\boldsymbol{H}^{(2)}$ are equinumerous results in
+$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}})+\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}})-2\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}}).$ (27)
 Note that $D$ is indeed a pseudometric since $k$ is positive definite as noted by Joshi et al.
@@
 By the reproducing property of $k$ and the linearity of the inner product, it may be shown that
-$\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert=D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right),$ (28)
+$\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert=D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right),$ (28)
 where
-$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h})$ (29)
+$G(\boldsymbol{H})=\sum_{\boldsymbol{h}\in\boldsymbol{H}}g(\boldsymbol{h})$ (29)
 and $g$ is the corresponding feature map of the kernel $k$. Hence, for every $\varepsilon_{1}>\varepsilon_{2}>0$,
-$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
-Appendix B Experimental set-up
-All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend.
+$\varepsilon_{2}<\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
+Appendix B Additional benchmark experiments
+Heterophilous Datasets
+Heterophilous Datasets (Platonov et al., 2023b) is a collection of node property prediction benchmark datasets for evaluating GNNs under heterophily.
+In particular, the roman-empire, amazon-ratings, tolokers, and questions datasets have uncountable node features.
+Meanwhile, the minesweeper dataset has countable node features.
+The performance metric of roman-empire and amazon-ratings is accuracy while the performance metric of the remaining datasets is ROC-AUC.
+Platonov et al.
+(2023b) provides more information regarding the individual datasets.
+Table 5 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, and GAT across the five benchmark datasets with the experimental set-up closely following that of Platonov et al.
+(2023b).
+The test performance for the GraphTransformer (Shi et al., 2021) and heterophilous GNNs—H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2021), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), and JacobiConv (Wang & Zhang, 2022)—are also presented as additional baselines.
+Unsurprisingly, SIR-GCN performs poorly on amazon-ratings, tolokers, and questions as these datasets exhibit near-zero label informativeness (Platonov et al., 2023a) between key node labels and query node labels, which severely limits the ability of SIR-GCN to learn meaningful relationships between neighboring nodes.
+Despite this inherent challenge, the performance gap in ROC-AUC between SIR-GCN and the top-performing models on tolokers and questions remains minimal, reflecting its robustness.
+Conversely, SIR-GCN outperforms classical GNNs, GraphTransformer, and all heterophilous GNNs on roman-empire and minesweeper, underscoring its utility even in heterophilous settings.
+Overall, the results highlight that SIR-GCN remains competitive in heterophilous graph tasks.
+Future works may further explore its theoretical and empirical properties under heterophily.
+Table 5: Test performance on Heterophilous Datasets.
+Model roman-empire ($\uparrow$) amazon-ratings ($\uparrow$) minesweeper ($\uparrow$) tolokers ($\uparrow$) questions ($\uparrow$)
+GCN 73.69 $\pm$ 0.74 48.70 $\pm$ 0.63 89.75 $\pm$ 0.52 83.64 $\pm$ 0.67 76.09 $\pm$ 1.27
+GraphSAGE 85.74 $\pm$ 0.67 53.63 $\pm$ 0.39 93.51 $\pm$ 0.57 82.43 $\pm$ 0.44 76.44 $\pm$ 0.62
+GAT 80.87 $\pm$ 0.30 49.09 $\pm$ 0.63 92.01 $\pm$ 0.68 83.70 $\pm$ 0.47 77.43 $\pm$ 1.20
+GraphTransformer 86.51 $\pm$ 0.73 51.17 $\pm$ 0.66 91.85 $\pm$ 0.76 83.23 $\pm$ 0.64 77.95 $\pm$ 0.68
+H2GCN 60.11 $\pm$ 0.52 36.47 $\pm$ 0.23 89.71 $\pm$ 0.31 73.35 $\pm$ 1.01 63.59 $\pm$ 1.46
+CPGNN 63.96 $\pm$ 0.62 39.79 $\pm$ 0.77 52.03 $\pm$ 5.46 73.36 $\pm$ 1.01 65.96 $\pm$ 1.95
+GPR-GNN 64.85 $\pm$ 0.27 44.88 $\pm$ 0.34 86.24 $\pm$ 0.61 72.94 $\pm$ 0.97 55.48 $\pm$ 0.91
+FSGNN 79.92 $\pm$ 0.56 52.74 $\pm$ 0.83 90.08 $\pm$ 0.70 82.76 $\pm$ 0.61 78.86 $\pm$ 0.92
+GloGNN 59.63 $\pm$ 0.69 36.89 $\pm$ 0.14 51.08 $\pm$ 1.23 73.39 $\pm$ 1.17 65.74 $\pm$ 1.19
+FAGCN 65.22 $\pm$ 0.56 44.12 $\pm$ 0.30 88.17 $\pm$ 0.73 77.75 $\pm$ 1.05 77.24 $\pm$ 1.26
+GBK-GNN 74.57 $\pm$ 0.47 45.98 $\pm$ 0.71 90.85 $\pm$ 0.58 81.01 $\pm$ 0.67 74.47 $\pm$ 0.86
+JacobiConv 71.14 $\pm$ 0.42 43.55 $\pm$ 0.48 89.66 $\pm$ 0.40 68.66 $\pm$ 0.65 73.88 $\pm$ 1.16
+SIR-GCN 87.67 $\pm$ 0.28 46.73 $\pm$ 0.61 94.12 $\pm$ 0.42 82.85 $\pm$ 0.72 75.33 $\pm$ 1.34
+Appendix C Experimental set-up
+All experiments are performed on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (DGL) (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend.
 For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values.
 For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials.
-The hyperparameters are chosen based on previous results and heuristics without extensive tuning.
 The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
-B.1 Synthetic datasets
+C.1 Synthetic datasets
 DictionaryLookup
 Adopting Brody et al.
-(2021), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration.
+(2022), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and/or values, while the test dataset comprises 1,000 bipartite graphs with the same configuration.
 All models utilize a single GNN layer with $4n$ hidden units.
-A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and std aggregators.
-Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
-GraphHeterophily
-The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration.
+A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and standard deviation aggregators.
+Model training is performed with the AdamW (Loshchilov & Hutter, 2019) optimizer for 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+HeteroEdgeCount
+The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with uniformly selected edges using the rand_graph function of DGL and uniformly assigned node labels from one of $c$ classes using the randint function of PyTorch.
+These measures ensure that the graphs are sufficiently diverse with respect to graph structure and heterophily.
+Meanwhile, the test dataset comprises 1,000 directed graphs with the same configuration.
 All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function.
-A feed-forward neural network is also used for GIN while PNA uses the sum, max, and std aggregators.
-Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
-B.2 Benchmark datasets
+A feed-forward neural network is also used for GIN while PNA uses the sum, max, and standard deviation aggregators.
+Model training is performed with the AdamW optimizer for 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+C.2 Benchmark datasets
 Benchmarking GNNs
@@
 In line with Dwivedi et al.
-(2023), all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000.
+(2023), Corso et al.
+(2020), and Tailor et al.
+(2022), all models utilize 4 GNN layers with batch normalization and residual connections while constrained to a parameter budget of 100,000.
 Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting.
@@
 Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10.
-Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
-The reported results for other models in Table 3 are obtained from Dwivedi et al.
+Model training is performed with the AdamW optimizer for a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
+The reported results for the other models in Table 3 are obtained from Dwivedi et al.
 (2023), Corso et al.
@@
 (2022).
-ogbn-arxiv
-The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al.
+Open Graph Benchmark
+The datasets are obtained from ogb, the Open Graph Benchmark Python package, with data splits following Hu et al.
+(2020a).
+In line with Corso et al.
+(2020) and Tailor et al.
+(2022), the model for ogbn-arxiv and ogbg-molhiv utilizes 3 and 4 GNN layers, respectively, with batch normalization and residual connections while constrained to a parameter budget of 100,000.
+Weight decays in $\left\{1\times 10^{-4},1\times 10^{-3}\right\}$ and dropouts with a rate of 0.2 are also used to prevent overfitting.
+The symmetric mean and max aggregators are used with mean pooling as the graph readout function for ogbg-molhiv.
+Model training is performed with the AdamW optimizer for a maximum of 1000 and 100 epochs, respectively, with a batch size of 64 for ogbg-molhiv and learning rates in $\left\{0.001,0.01\right\}$ that decays by a factor of 0.5 with patience of 40 and 10 epochs, respectively, based on the training loss.
+The reported results for the other models in Table 4 are obtained from Corso et al.
+(2020) and Tailor et al.
 (2022).
-The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections.
-Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting.
-The symmetric mean aggregator is used along with existing tricks in literature.
-Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss.
-The reported results for other models in Table 4 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
-ogbg-molhiv
-The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al.
-(2020).
-The models utilize a single GNN layer, modified to leverage edge features as described in Appendix D, with 300 hidden units, batch/graph normalization, and residual connections.
-Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting.
-The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function.
-Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss.
-The reported results for other models in Table 5 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
-Appendix C Model run time
-As an additional evaluation, the run time per epoch for each model in the synthetic datasets is presented in Tables 6 and 7.
-The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity.
-Notably, SIR-GCN is computationally less expensive than PNA while delivering superior performance.
-These results further underscore the practicality of the proposed model.
-Table 6: Run time per epoch on DictionaryLookup.
+Heterophilous Datasets
+The datasets are obtained from dgl with data splits following Platonov et al.
+(2023b).
+In line with Platonov et al.
+(2023b), the number of GNN layers is chosen from $\left\{3,5\right\}$, employing residual connections and dropouts with a rate of 0.2.
+Moreover, the hidden dimension is chosen from $\left\{256,512\right\}$, utilizing batch and layer normalization as well as the symmetric mean and max aggregators.
+Model training is performed with the AdamW optimizer for a maximum of 1000 epochs and a learning rate of $3\times 10^{-5}$.
+The reported results for the other models in Table 5 are obtained from Platonov et al.
+(2023b).
+Appendix D Runtime analysis
+As an additional evaluation, the inference runtime for each model in the synthetic datasets is presented in Tables 6 and 7.
+The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity, specifically with regards to PNA which is also designed for uncountable node features.
+Table 6: DictionaryLookup inference runtime.
 Model $n=10$ $n=20$ $n=30$ $n=40$ $n=50$
@@
 SIR-GCN 0.4687s $\pm$ 0.0777s 0.6066s $\pm$ 0.0398s 0.8053s $\pm$ 0.0485s 1.1496s $\pm$ 0.0427s 1.7031s $\pm$ 0.0458s
-Table 7: Run time per epoch on GraphHeterophily.
+Table 7: HeteroEdgeCount inference runtime.
 Model $c=2$ $c=4$ $c=6$ $c=8$ $c=10$
@@
 SIR-GCN 0.5338s $\pm$ 0.0353s 0.5264s $\pm$ 0.0737s 0.5635s $\pm$ 0.0695s 0.5764s $\pm$ 0.0401s 0.6230s $\pm$ 0.0388s
-Appendix D SIR-GCN extensions
-Denote $\bm{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$.
-Following the intuition presented in Eqns.
-8 and 9, SIR-GCN with residual connection may be modified to leverage edge features to obtain
-$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{E}}\bm{h_{u,v}}+\bm{W_{K}}\bm{h_{v}}\right),$ (31)
-where $\bm{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
+Table 8 further complements these results by presenting the asymptotic computational runtime complexity of the different GNNs.
+In particular, SIR-GCN first computes the linear transformations $\boldsymbol{W_{Q}}\boldsymbol{h_{u}}$ and $\boldsymbol{W_{K}}\boldsymbol{h_{v}}$ for every node which incurs $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}\right)$.
+Afterward, $\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right)$ is computed for every edge, using the previously calculated values, and aggregated across the neighbors of each node which incurs $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\right)$.
+Finally, the aggregated values are linearly transformed with $\boldsymbol{W_{R}}$ for every node which incurs $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ [^4] In the case of SIR-GCN with max aggregation, the linear transformation and the max aggregator cannot be interchanged.
+Hence, the linear transformation $\boldsymbol{W_{R}}$ must be performed along edges which incurs $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$..
+In total, since SIR-GCN employs only linear transformations along nodes and only an activation function along edges, its computational complexity is comparable to GCN, GraphSAGE, GAT, GATv2, and GIN.
+Specifically, these models achieve computational efficiency by maintaining linear complexity along edges attributed to activation functions and neighborhood aggregation.
+Despite this, SIR-GCN consistently outperforms these classical GNNs across all benchmarks.
+Notably, SIR-GCN also demonstrates a lower complexity than PNA due to the number of aggregators used, yet delivers superior performance across most datasets.
+These additional analyses further underscore the practical utility of the proposed model.
+Table 8: Asymptotic runtime complexity.
+Model Complexity
+GCN $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
+GraphSAGE $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
+GAT/GATv2 $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
+GIN $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{in}}+\left|\mathcal{V}\right|\times\texttt{MLP}\right)$
+PNA $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{in}}^{2}+\left|\mathcal{E}\right|\times d_{\text{in}}\times k+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}\times k\right)$
+SIR-GCN $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$
+Note: $k$ represents the number of aggregators and scalers in PNA.
+Appendix E SIR-GCN extensions
+Denote $\boldsymbol{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$.
+Following the intuition presented in Section 3.1, SIR-GCN with residual connection may be modified to leverage edge features to obtain
+$\boldsymbol{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\boldsymbol{h_{u}})+\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{E}}\boldsymbol{h_{u,v}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (31)
+where $\boldsymbol{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
 $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times\texttt{MLP}_{\text{Res}}\right),$ (32)
-with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA.
-Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2019).
-Furthermore, one may inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$.
-For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2021) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.
+where $\texttt{MLP}_{\text{Res}}$ denotes the computational complexity of $\textsc{MLP}_{\text{Res}}$, making it comparable to PNA.
+Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2020b).
+Furthermore, one may also inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$.
+For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2023) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.